### v8

Abstract
We describe a new dependent-rounding algorithmic framework for bipartite graphs. Given a fractional assignment $\vec{x}$ of values to edges of a graph $G=(U\cup V,E)$, the algorithms return an integral solution $\vec{X}$ such that each right-node $v\in V$ has at most one neighboring edge $f$ with $X_{f}=1$, and the variables $X_{e}$ also satisfy broad nonpositive-correlation properties. In particular, for any edges $e_{1},e_{2}$ sharing a left-node $u\in U$, the variables $X_{e_{1}},X_{e_{2}}$ have strong negative correlation, i.e. the expectation of $X_{e_{1}}X_{e_{2}}$ is significantly below $x_{e_{1}}x_{e_{2}}$.
This algorithm is based on generating negatively-correlated Exponential random variables and using them for a rounding method inspired by a contention-resolution scheme of Im & Shadloo (2020). Our algorithm gives stronger and much more flexible negative correlation properties.
Dependent rounding schemes with negative correlation properties have been used for approximation algorithms for job-scheduling on unrelated machines to minimize weighted completion times (Bansal, Srinivasan, & Svensson (2021), Im & Shadloo (2020), Im & Li (2023)). Using our new dependent-rounding algorithm, among other improvements, we obtain a $1.398$-approximation for this problem. This significantly improves over the prior $1.45$-approximation ratio of Im & Li (2023).
1 Introduction
Many discrete optimization algorithms are based on the following framework: we start by solving some relaxation of the problem instance (e.g., a linear program), obtaining a fractional solution $\vec{x}=(x_{1},\dots,x_{n})$. We then want to convert it into an integral solution $\vec{X}=(X_{1},\dots,X_{n})$, with probabilistic properties related to $\vec{x}$, while also satisfying any needed hard combinatorial constraints for the problem. This general framework often goes by the name dependent rounding, since in general the combinatorial constraints will necessarily induce dependencies among the variables $X_{i}$.
Some forms of dependent rounding are highly tailored to specific algorithmic problems, while others are very general. These frequently appear in clustering problems, for instance, where there is a hard constraint that every data item must be mapped to a cluster-center; see for example [7, 4]. It also appears in a number of job-scheduling problems [6], where there is a hard constraint that every job must be assigned to some processing node.
One particularly important property is that the variables $X_{i}$ should satisfy nonpositive-correlation properties; namely, for certain subsets $L\subseteq\{1,\dots,n\}$, we should have
${\mathbb{E}}\Bigl[\prod_{i\in L}X_{i}\Bigr]\leq\prod_{i\in L}{\mathbb{E}}[X_{i}]=\prod_{i\in L}x_{i}$ (1)
For example, this property leads to Chernoff-type concentration bounds. Depending on the combinatorial constraints, such an inequality may only be satisfied for certain limited choices of subset $L$. Note that independent rounding would certainly satisfy this property for all sets $L$.
For some combinatorial problems, nonpositive-correlation is not enough; we need strong negative correlation. Namely, for certain edge sets $L$, we want a bound of the form
${\mathbb{E}}\Bigl[\prod_{i\in L}X_{i}\Bigr]\leq(1-\phi)\cdot\prod_{i\in L}x_{i}$ (2)
for some parameter $\phi\gg 0$; the precise value of $\phi$ may depend on the set $L$ and/or the values $x_{i}$.
One particularly interesting application comes from job scheduling on unrelated machines to minimize completion time. We will discuss this problem in more detail later, but for now let us provide a brief summary. We have a set of machines $\mathcal{M}$ and a set of jobs $\mathcal{J}$, where each job $j\in\mathcal{J}$ has a given processing time on each machine $i\in\mathcal{M}$. We want to minimize the objective function $\sum_{j\in\mathcal{J}}w_{j}C_{j}$, where $C_{j}$ is the completion time of $j$ on its chosen machine and $w_{j}$ is a weight function. A long series of approximation algorithms have been developed for this problem, often based on sophisticated convex relaxations and dependent-rounding algorithms.
The aim of this paper is to obtain a more unified and general picture of dependent rounding with strong negative correlation. Our algorithm, along with some other improvements, leads to a $1.398$-approximation algorithm for the above machine-scheduling problem, which improves over the algorithm of [10] while also being simpler and more generic.
1.1 Definitions and results for the bipartite-rounding setting
Many combinatorial assignment problems, such as machine-scheduling, can be abstracted in the following framework: we consider a bipartite graph $G=(U\cup V,E)$, where the fractional variables $\vec{x}\in[0,1]^{E}$ are associated to its edges. Our goal is to generate rounded random variables $\vec{X}\in\{0,1\}^{E}$ which match the fractional variables, while also achieving nonpositive-correlation or strong negative correlation among certain subset of values $X_{e}$ corresponding to the vertices.
We call $U$ and $V$ left-nodes and right-nodes respectively. For any vertex $w\in U\cup V$ we define $\Gamma(w)$ to be the edges incident to $w$.
The bipartite rounding algorithm of [6] is a powerful example of this framework. This algorithm, which has found numerous applications in optimization problems, guaranteed a limited case of Eq. (1): namely, it held for any edge set $L$ which was a subset of the neighborhood of any vertex.
Our first main contribution, in Sections 2 and 3, is to develop a new bipartite-graph rounding algorithm based on negatively correlated Exponential random variables. The construction has two stages. First, we give a general method of generating Exponential random variables with certain types of negative correlation. Second, we use this in a graph contention-resolution scheme, which takes as input a “rate” vector $\vec{\rho}\in[0,1]^{E}$; we will discuss its role shortly.
Let us say that a set of edges $L\subseteq E$ is stable if it does not contain any pair of edges of the form $e=(u,v),e^{\prime}=(u^{\prime},v^{\prime})$, where $(u,v^{\prime})$ is also an edge in $E$. Equivalently, $L$ does not contain any pair of edges whose distance is precisely two in the line graph of $G$.
We get the following result:
Theorem 1.
Suppose $\sum_{e\in\Gamma(v)}x_{v}\leq 1$ for all right-nodes $v\in V$, and $\sum_{e\in\Gamma(u)}\rho_{u}\leq 1$ for all left-nodes $u\in U$. Then the algorithm DepRound satisfies the following properties (A1) — (A4):
  • (A1)
    For each right-node $v\in V$, we have $\sum_{e\in\Gamma(v)}X_{e}\leq 1$ with probability one.
  • (A2)
    For each edge $e$ there holds ${\mathbb{E}}[X_{e}]=x_{e}$
  • (A3)
    For any stable edge-set $L\subseteq E$, there holds ${\mathbb{E}}[\prod_{e\in L}X_{e}]\leq\prod_{e\in L}x_{e}$
  • (A4)
    For any pair of edges $e_{1}=(u,v_{1}),e_{2}=(u,v_{2})$ sharing a common left-node, we have
    ${\mathbb{E}}[X_{e_{1}}X_{e_{2}}]\leq(1-\Phi(x_{e_{1}},x_{e_{2}};\rho_{e_{1}},\rho_{e_{2}}))x_{e_{1}}x_{e_{2}},$
    where we define the function $\Phi(x_{1},x_{2};\rho_{1},\rho_{2}):[0,1]^{4}\rightarrow[0,1]$ by
    $\Phi(x_{1},x_{2};\rho_{1},\rho_{2})=\begin{cases}\frac{((1-\rho_{1})^{1-1/x_{1}}-1)((1-\rho_{2})^{1-1/x_{2}}-1)}{(1-\rho_{1})^{1-1/x_{1}}(1-\rho_{2})^{1-1/x_{2}}+\rho_{1}+\rho_{2}-1}&\text{for $x_{1},x_{2},\rho_{1},\rho_{2}\in(0,1)$}\\ 0&\text{otherwise}\end{cases}$
There is the question of choosing the parameter $\rho$. This vector $\vec{\rho}$ is not necessarily related to $\vec{x}$; it can be carefully chosen to obtain different “shapes” of negative correlation. Intuitively, $\rho_{e}$ controls how much anti-correlation $e$ should have with other edges. We emphasize that there is no single optimal choice, and our scheduling algorithm will use this flexibility in a somewhat subtle way. However, there are a few natural parameterizations, as given in the following result:
Corollary 2.
For a left-node $u$, let $\lambda_{u}=\sum_{e\in\Gamma(u)}x_{e}$. If we set $\rho_{e}=1-\mathrm{e}^{-x_{e}/\lambda_{u}}$ for all $e\in\Gamma(u)$, then for any edges $e_{1},e_{2}$ with $x_{e_{1}}x_{e_{2}}>0$, we have [^3] Here and throughout we write $\mathrm{e}=2.718$... We use a different font to distinguish the constant $\mathrm{e}$ from an edge $e$ in a graph.
$\Phi(x_{e_{1}},x_{e_{2}};\rho_{e_{1}},\rho_{e_{2}})\geq 1-\frac{\mathrm{e}^{x_{e_{1}}/\lambda_{u}}+\mathrm{e}^{x_{e_{2}}/\lambda_{u}}}{1+\mathrm{e}^{1/\lambda_{u}}}$
Moreover, if $\lambda_{u}\leq 3/4$ and we set $\rho_{e}=x_{e}/\lambda_{u}$ for all $e\in\Gamma(u)$, then
$\Phi(x_{e_{1}},x_{e_{2}};\rho_{e_{1}},\rho_{e_{2}})\geq\frac{\mathrm{e}^{1/\lambda_{u}}-1}{\mathrm{e}^{1/\lambda_{u}}+1}-0.57(x_{e_{1}}+x_{e_{2}})\max\{0,\lambda_{u}-0.45\}$
The first result strictly generalizes the result of [9] (which only considered the setting with $\lambda_{u}=1$). The second result can be stronger quantitatively, and also has a simpler algebraic formula.
We note that there is a closely related setting, as originally considered in [2, 3, 9], where we have a bipartite graph $G$ where the edges incident to each left-node are partitioned into “blocks.” In this setting, we should have nonpositive-correlation among edges incident to each left-node, and strong negative correlation within each block. Our framework can also handle this setting, by transforming each block of the graph $G$ to a new separate left-node. Note that the neighborhood of a vertex in the original graph $G$ will form a stable edge-set in the resulting graph $G^{\prime}$.
We emphasize that the dependent rounding algorithm, and its analysis, is completely self-contained and does not depend on the machine-scheduling setting. Because of its generality, it may be applicable to other combinatorial optimization problems.
1.2 Machine-scheduling on unrelated machines
This problem is denoted $R||\sum_{j}w_{j}C_{j}$ in the common nomenclature for machine scheduling. Here, we have a set of machines $\mathcal{M}$ and set of jobs $\mathcal{J}$, where each job $j$ has a weight $w_{j}$ and has a separate processing time $p^{(i)}_{j}$ on each machine $i$. Our goal is to assign the jobs to the machines in order, so as to minimize the overall weighted completion time $\sum_{j}w_{j}C_{j}$, where $C_{j}$ is the sum of $p^{(i)}_{j^{\prime}}$ over all jobs assigned on machine $i$ up to and including $j$.
On a single machine, there is a simple greedy heuristic for this problem: jobs should be scheduled in non-increasing order of the ratio $\sigma_{j}=w_{j}/p_{j}$, which is known as the Smith ratio [18]. For multiple machines, it is an intriguing and long-studied APX-hard problem [8]. It has attracted attention, in part, because it leads to sophisticated convex programming relaxations and rounding algorithms. Since the 2000’s, there were a series of $1.5$-approximation algorithms based on various non-trivial fractional relaxations [15, 17, 16].
As shown in [2], going beyond this approximation ratio demands much more involved algorithms: some of the main convex relaxations have integrality gap $1.5$, and furthermore rounding strategies which treat each job independently, as had been used in all previous algorithms, are inherently limited to approximation ratio $1.5$. In a breakthrough, [2] achieved a $1.5-\varepsilon$ approximation factor for some minuscule constant $\varepsilon>0$. There have been further improvements [12, 9, 10]; most recently before this work, [10] gave a $1.45$-approximation (see also [13] for a faster implementation of that algorithm).
These newer algorithms can all be described in the same general framework. First, they solve an appropriate relaxation, giving fractional assignments $x^{(i)}_{j}$. Second, they group the jobs on each machine, forming “clusters” of jobs with similar processing times. Finally, they apply some form of dependent rounding with strong negative correlation properties within each cluster to convert this into an integral assignment.
There are two main types of relaxations for the first step. The work [2] used a semidefinite-programming (SDP) relaxation. The works [12, 9, 10] used a relaxation based on a time-indexed LP. It is not clear which of these relaxations is better, for example, is faster to solve, has a better approximation algorithm, or has a smaller integrality gap. We note that the time-indexed LP may be able to accommodate some other variants of the problem, such as having “release times” for each job, which do not seem possible for the SDP relaxation.
There have also been many different types of rounding algorithms used. The work [2] used a method based on a random walk in a polytope, which was also used as a black-box subroutine by [12]. This rounding method was subsequently improved in [3]. The work [9] developed a very different rounding algorithm based on contention resolution of Poisson processes. Finally, [10] used a specialized rounding algorithm closely tied to the time-indexed LP structure; it borrows some features from both the random-walk and contention-resolution algorithms.
Our new algorithm will also fit into this framework, where we use the SDP as a starting point and we use our algorithm DepRound for the rounding. We show the following:
Theorem 3.
There is a randomized approximation algorithm for Scheduling on Unrelated Machines to Minimize Weighted Completion Time with approximation ratio $1.398$.
In particular, the SDP relaxation has integrality gap at most $1.398$.
This is the first improvement to the integrality gap for the SDP relaxation since the original work [2]; all later improvements have been based on the time-indexed LP. It is interesting now to determine which fractional relaxation has a better gap.
To explain our improvement, we note that it is relatively straightforward to handle a scenario with many jobs of small mass and similar processing times. The difficulty lies in grouping disparate items together. This typically results in “ragged” clusters and, what is worse, the “leftover clusters” for each processing-time class. A large part of our improvement is technical, coming from tracking the contributions of leftover clusters more carefully. There are two algorithmic ideas to highlight.
First, the DepRound algorithm is much more flexible for rounding: its negative correlation guarantees scale with the size of each cluster or even the individual items, as opposed to the maximum cluster size. The parameter $\rho$ available in DepRound allows us to put less “anti-correlation strength” on the final job within a cluster. This makes the clusters act in a significantly more uniform way irrespective of their total mass. In particular, we no longer need to deal separately with jobs which have “large” mass on a given machine.
Second, we use a random shift before quantizing the items by processing time, which ensures that items are more evenly distributed within each class. This technique was also used in [9, 10], but had not been analyzed in the context of the SDP relaxation.
Note:
After the conference version of this paper, [14] proposed a new algorithm for this problem with approximation ratio $1.36$. It uses different techniques, based on iterated rounding. Its approximation ratio (and integrality gap) are shown in terms of the Configuration LP, which is stronger than either the time-indexed LP or SDP relaxations.
2 Negatively-correlated Exponential random variables
Before we consider bipartite rounding, we derive a more basic result in probability theory: how to generate Exponential random variables with strong negative correlations. The analysis relies heavily on properties of negatively-associated (NA) random variables [11]. Formally, we say that random variables $X_{1},\dots,X_{k}$ are NA if for any disjoint subsets $A,B\subseteq\{1,...,k\}$ and any increasing functions $f,g$, there holds
${\mathbb{E}}[f(X_{i}:i\in A)g(X_{j}:j\in B)]\leq{\mathbb{E}}[f(X_{i}:i\in A)]{\mathbb{E}}[g(X_{j}:j\in B)]$ (3)
We quote a few useful facts about such variables from [11, 19].
Theorem 4.
  1. 1.
    If $X_{1},\dots,X_{k}$ are NA, then ${\mathbb{E}}[X_{1}\cdots X_{k}]\leq{\mathbb{E}}[X_{1}]\cdots{\mathbb{E}}[X_{k}]$.
  2. 2.
    If $X_{1},\dots,X_{k}$ are zero-one random variables with $X_{1}+\dots+X_{k}\leq 1$, then $X_{1},\dots,X_{k}$ are NA.
  3. 3.
    If $\mathcal{X},\mathcal{X}^{\prime}$ are collections of NA random variables, and the joint distribution of $\mathcal{X}$ is independent from that $\mathcal{X}^{\prime}$, then $\mathcal{X}\cup\mathcal{X}^{\prime}$ are NA.
  4. 4.
    If $X_{1},\dots,X_{k}$ are NA random variables and $f_{1},\dots,f_{\ell}$ are functions defined on disjoint subsets of $\{1,\dots,k\}$, such that all $f_{1},\dots,f_{\ell}$ are monotonically non-increasing or all $f_{1},\dots,f_{\ell}$ are monotonically non-decreasing, then random variables $f_{i}(\vec{X}):i=1,\dots,\ell$ are NA. In particular, ${\mathbb{E}}[f_{1}(\vec{X})\cdots f_{\ell}(\vec{X})]\leq{\mathbb{E}}[f_{1}(\vec{X})]\cdots{\mathbb{E}}[f_{\ell}(\vec{X})]$.
Our algorithm also makes use of the Multivariate Geometric distribution [1]: given a probability vector $\vec{p}\in[0,1]^{n}$, this distribution on vector $(X_{1},\dots,X_{n})$ is defined by running an infinite sequence of experiments, where each $j^{\text{th}}$ experiment takes on value $i$ with probability $p_{i}$, and setting $X_{i}$ to be the number of trials before first seeing $i$. We note a few simple properties of this distribution.
Observation 5.
Let $(X_{1},\dots,X_{n})$ be a Multivariate Geometric random variable with rate $\vec{p}$. Then the variables $X_{1},\dots,X_{n}$ are NA and each marginal distribution $X_{i}$ is Geometric with rate $p_{i}$.
Proof.
Observe that $X_{1},\dots,X_{n}$ can be viewed in terms of the following infinite sequence of random variables: for each $j\in\mathbb{Z}_{\geq 0}$, draw variable $F_{j}\in\{1,\dots,n\}$ with $\Pr(F_{j}=i)=p_{i}$ for all $i,j$. Then set $X_{i}=\min\{j:F_{j}=i\}$ for each $i$. Let $B_{i,j}$ be the indicator random variable for the event $F_{j}=i$. For each $j$, the values $B_{i,j}$ are zero-one random variables which sum to one, so they are NA. Since there is no interaction between the different indices $j$, all random variables $B_{i,j}$ are NA. Furthermore, each $X_{i}$ is a monotone-down function of the values $B_{i,j}$. Hence, all variables $X_{i}$ are NA.
Furthermore, each $X_{i}$ is the waiting time until we see the first variable $B_{i,j}$; since these are Bernoulli with rate $p_{i}$, the distribution of $X_{i}$ is Geometric with rate $p_{i}$
We describe our algorithm to generate correlated unit Exponential random variables $Z_{1},\dots,Z_{n}$. It takes as input a vector $\vec{\rho}\in[0,1]^{n}$ with $\sum_{i}\rho_{i}\leq 1$, which will determine the covariances.
1Set $\rho_{0}=1-(\rho_{1}+\dots+\rho_{n})$
2 Set $(X_{0},\dots,X_{n})\leftarrow\textsc{MultivariateGeometric}(\vec{\rho})$.
3 for $i=1,\dots,n$ do
4 if $\rho_{i}\in(0,1)$ then
5      Set $\alpha_{i}=-\log(1-\rho_{i})$
6      Draw random variable $S_{i}\in[0,1]$ with probability density function $\alpha_{i}\mathrm{e}^{-\alpha_{i}s}/\rho_{i}$
7      Set $Z_{i}=\alpha_{i}(X_{i}+S_{i})$
8
9 else
10      Draw $Z_{i}$ as an independent Exponential random variable with rate 11
11
return vector $Z$
Algorithm 1 $\textsc{CorrelatedExponential}(\vec{\rho})$
The density function at Line 5 is valid in that $\int_{s=0}^{1}\alpha_{i}\mathrm{e}^{-\alpha_{i}s}/\rho_{i}\ \mathrm{d}s=\frac{1}{\rho_{i}}(1-\mathrm{e}^{-\alpha_{i}})=1$. The algorithm can be implemented to run in randomized polynomial time via standard sampling procedures. We now analyze its probabilistic properties.
Proposition 6.
Each random variable $Z_{i}$ has an Exponential distribution with rate $1$, i.e. it has probability density function $e^{-z}:z\in[0,\infty)$.
Proof.
It is clear if $\rho_{i}\in\{0,1\}$, so suppose $\rho_{i}\in(0,1)$ and consider random variable $L=X_{i}+S_{i}$. We claim that $L$ has an Exponential distribution with rate $\alpha_{i}$. For, consider some $z\in\mathbb{R}_{\geq 0}$, where $z=j+s$ and $j\in\mathbb{Z}_{\geq 0},s\in[0,1)$. Since $X_{i}$ is Geometric with rate $\rho_{i}$, the probability density function for $L$ at $z$ is given by
$\rho_{i}(1-\rho_{i})^{j}\cdot\frac{\alpha_{i}\mathrm{e}^{-\alpha_{i}s}}{\rho_{i}}=(\mathrm{e}^{-\alpha_{i}})^{j}\cdot\alpha_{i}\mathrm{e}^{-\alpha_{i}s}=\mathrm{e}^{-\alpha_{i}(j+s)}\cdot\alpha_{i}=\alpha_{i}\cdot\mathrm{e}^{-\alpha_{i}z}$
which is precisely the pdf of a rate-$\alpha_{i}$ Exponential random variable. Since rescaling Exponential random variables changes their rate, the variable $Z_{i}$ is Exponential with rate $1$. ∎
Proposition 7.
The random variables $Z_{1},\dots,Z_{n}$ are NA.
Proof.
The random variables $X_{i}$ are NA by Observation 5. The $S_{i}$ variables are clearly NA since they are independent. Each variable $Z_{i}$ for $\rho_{i}\in(0,1)$ is an increasing function of the random variables $S_{i},X_{i}$. Hence, these $Z$ variables are also NA. The $Z_{i}$ variables with $\rho_{i}\in\{0,1\}$ are NA since they independent of all other variables. ∎
The crucial property is that the variables $Z_{i}$ have significant negative correlation (depending on shape parameter $\vec{\rho}$). Specifically, we have the following:
Lemma 8.
For any indices $i_{1},i_{2}$ with $\rho_{i_{1}},\rho_{i_{2}}\in(0,1)$, and any values $q_{1},q_{2}\in(-\infty,1)$, we have
${\mathbb{E}}[\mathrm{e}^{q_{1}Z_{i_{1}}+q_{2}Z_{i_{2}}}]=\frac{1}{(1-q_{1})(1-q_{2})}\cdot\Bigl(1-\frac{((1-\rho_{i_{1}})^{q_{1}}-1)((1-\rho_{i_{2}})^{q_{2}}-1)}{(1-\rho_{i_{1}})^{q_{1}}(1-\rho_{i_{2}})^{q_{2}}+\rho_{i_{1}}+\rho_{i_{2}}-1}\Bigr)$
Proof.
Suppose without loss of generality that $i_{1}=1,i_{2}=2$. For $i=1,2$ define parameter $\theta_{i}=\alpha_{i}q_{i}=-q_{i}\log(1-\rho_{i})$ and random variable $L_{i}=X_{i}+S_{i}$. We calculate:
${\mathbb{E}}[\mathrm{e}^{q_{1}Z_{1}+q_{2}Z_{2}}]={\mathbb{E}}[\mathrm{e}^{\theta_{1}L_{1}+\theta_{2}L_{2}}]={\mathbb{E}}[\mathrm{e}^{\theta_{1}S_{1}}]{\mathbb{E}}[\mathrm{e}^{\theta_{2}S_{2}}]{\mathbb{E}}[\mathrm{e}^{\theta_{1}X_{1}+\theta_{2}X_{2}}]$
where the last equality holds since the variables $S_{1},S_{2}$ are independent of $X$.
We will calculate these term by term. For $i=1,2$ we have:
${\mathbb{E}}[\mathrm{e}^{\theta_{i}S_{i}}]$ $=\frac{\alpha_{i}}{\rho_{i}}\int_{s=0}^{1}\mathrm{e}^{-\alpha_{i}s+\theta_{i}s}\ \mathrm{d}s=\frac{\alpha_{i}(1-\mathrm{e}^{\theta_{i}-\alpha_{i}})}{\rho_{i}(\alpha_{i}-\theta_{i})}$
We calculate the joint distribution of $X_{1},X_{2}$ as:
$\Pr(X_{1}=x_{1},X_{2}=x_{2})=\begin{cases}(1-\rho_{1}-\rho_{2})^{x_{1}}\rho_{1}(1-\rho_{2})^{x_{2}-x_{1}-1}\rho_{2}&\text{for $0\leq x_{1}<x_{2}$}\\ (1-\rho_{1}-\rho_{2})^{x_{2}}\rho_{2}(1-\rho_{1})^{x_{1}-x_{2}-1}\rho_{1}&\text{for $0\leq x_{2}<x_{1}$}\\ 0&\text{otherwise}\end{cases}$
So we can sum over pairs $x_{1},x_{2}$ to get:
${\mathbb{E}}[\mathrm{e}^{\theta_{1}X_{1}+\theta_{2}X_{2}}]$ $=\sum_{0\leq x_{1}<x_{2}}(1-\rho_{1}-\rho_{2})^{x_{1}}\rho_{1}(1-\rho_{2})^{x_{2}-x_{1}-1}\rho_{2}\mathrm{e}^{\theta_{1}X_{1}+\theta_{2}X_{2}}$
$\qquad\qquad+\sum_{0\leq x_{2}<x_{1}}(1-\rho_{1}-\rho_{2})^{x_{2}}\rho_{2}(1-\rho_{1})^{x_{1}-x_{2}-1}\rho_{1}\mathrm{e}^{\theta_{1}X_{1}+\theta_{2}X_{2}}$
$=\frac{\rho_{1}\rho_{2}}{1-\mathrm{e}^{\theta_{1}+\theta_{2}}(1-\rho_{1}-\rho_{2})}\cdot\Bigl(\frac{1}{\mathrm{e}^{-\theta_{1}}+\rho_{1}-1}+\frac{1}{\mathrm{e}^{-\theta_{2}}+\rho_{2}-1}\Bigr)$
The result then follows by multiplying the formulas for ${\mathbb{E}}[\mathrm{e}^{\theta_{1}X_{1}+\theta_{2}X_{2}}],{\mathbb{E}}[\mathrm{e}^{\theta_{1}S_{1}}],{\mathbb{E}}[\mathrm{e}^{\theta_{2}S_{2}}]$ and substituting for $\alpha_{i}=-\log(1-\rho_{i}),\theta_{i}=-q_{i}\log(1-\rho_{i})$. ∎
3 Bipartite dependent rounding algorithm
We consider here a variant of a rounding algorithm from [9], which was in turn inspired by an earlier “fair sharing” algorithm of [5]. For motivation, consider the following natural rounding procedure: each edge $e$ draws an Exponential random $Z_{e}$ variable with rate $x_{e}$. Then, for each right-node $v$, we set $X_{f}=1$ for the edge $f=\operatorname{argmin}_{e\in\Gamma(v)}X_{e}$. It is easy to see that this is equivalent to independent rounding, and in particular we have $\Pr(X_{e}=1)=x_{e}$.
Instead of generating the variables $Z_{e}$ independently, we will use our algorithm CorrelatedExponential. This creates the desired negative correlation in the bipartite rounding.
In describing our algorithm, we will assume that our original fraction solution $\vec{x}$ satisfies
$\forall e\in E\ \ x_{e}\in(0,1),\qquad\text{and}\qquad\forall v\in V\sum_{e\in\Gamma(v)}x_{e}=1$
These conditions can be assumed without loss of generality by removing edges with $x_{e}\in\{0,1\}$, and by adding dummy edges for each right-node.
1for each left-node $u$ do
2   Call $\vec{Z}^{(u)}\leftarrow\textsc{CorrelatedExponential}(\vec{\rho}^{(u)})$ for vector $(\rho^{u}(v)=\rho_{(u,v)}:v\in\Gamma(u))$
3Combine all vectors $\vec{Z}^{(u)}$ into a single vector $Z\in\mathbb{R}_{\geq 0}^{E}$ with $Z_{(u,v)}=Z^{(u)}_{v}$ for all $u,v$.
4 for each right-node $v$ do
5   Choose neighbor $e=\operatorname{argmin}_{f\in\Gamma(v)}Z_{f}/x_{f}$.
6   Set $X_{e}=1$ and $X_{e^{\prime}}=0$ for all other neighbors $e^{\prime}\in\Gamma(v)\setminus\{e\}$
7
return vector $\vec{X}$
Algorithm 2 $\textsc{DepRound}(\vec{\rho},\vec{x})$
From properties of CorrelatedExponential, the following properties immediately hold:
Proposition 9.
Suppose that $\vec{\rho}\in[0,1]^{E}$ satisfies $\sum_{e\in\Gamma(u)}\rho_{e}=1$ for all left-nodes $u$. Then DepRound can be implemented in polynomial time with the following properties:
  • Each variable $Z_{e}$ has the Exponential distribution with rate $1$.
  • The random variables $Z_{e}:e\in E$ are NA.
  • For a right-node $v$, all random variables $Z_{e}:e\in\Gamma(v)$ are independent.
For an edge-set $L$, let us write $Z_{L}$ for the vector of random variables $(Z_{e}:e\in L)$. At this point, we can immediately show property (A2):
Proposition 10.
For any edge $e$ there holds $\Pr(X_{e}=1)=x_{e}$.
Proof.
The random variables $Z_{f}:f\in\Gamma(v)$ are independent unit-rate Exponentials. So random variables $Z_{f}/x_{f}:f\in\Gamma(v)$ are independent Exponentials with rates $x_{f}$ respectively, and $\sum_{f\in\Gamma(v)}x_{f}=1$. It is a well-known standard fact that, for independent Exponential random variables $Y_{1},\dots,Y_{\ell}$ with rates $\lambda_{1},\dots,\lambda_{\ell}$, there holds $\Pr(Y_{i}=\min\{Y_{1},\dots,Y_{\ell}\})=\frac{\lambda_{i}}{\lambda_{1}+\dots+\lambda_{\ell}}$. ∎
We next turn to show properties (A3) and (A4). For this, we have the following key lemma:
Lemma 11.
Let $L\subseteq E$ be a stable edge-set. If we reveal the random variables $Z_{L}$, then we have
${\mathbb{E}}\Bigl[\prod_{e\in L}X_{e}\mid Z_{L}\Bigr]\leq\prod_{e\in L}\mathrm{e}^{(1-1/x_{e})\cdot Z_{e}}$
Proof.
We assume that all right-nodes of edges in $L$ are distinct, as otherwise $\prod_{e\in L}X_{e}=0$ with probability one.
Define $L^{\prime}$ to be the set of edges outside $L$ which share a right-node with an edge in $L$, that, is, the set of edges of the form $f=(u,v)\notin L$ where $(u^{\prime},v)\in L$. Let $W,W^{\prime}$ denote the set of left-nodes of edges of $L,L^{\prime}$ respectively. We claim that $W,W^{\prime}$ are disjoint. For, suppose $u\in W\cap W^{\prime}$. So there edges $(u,v)\in L,(u,v^{\prime})\in L^{\prime}$; by definition of $L^{\prime}$, there must be a corresponding edge $(u^{\prime},v^{\prime})\in L$. Since $G$ is a simple graph, necessarily $u\neq u^{\prime}$. Also $v\neq v^{\prime}$ since edges in $L$ have distinct right-nodes. This contradicts that $L$ is a stable set.
Suppose that we condition on all random variables corresponding to the nodes in $W$, in particular, we reveal all values $Z_{L}$. The random variables corresponding to nodes in $W^{\prime}$ have their original unconditioned probability distributions. We now have
$\Pr(\bigwedge_{e\in L}X_{e}=1)=\Pr\Bigl(\bigwedge_{(u,v)\in L}Z_{(u,v)}/x_{(u,v)}=\min_{f\in\Gamma(v)}Z_{f}/x_{f}\Bigr)$
where, here and in the remainder of the proof, we omit the conditioning on $W$ for brevity.
Each event $Z_{(u,v)}/x_{(u,v)}=\min_{f\in\Gamma(v)}Z_{f}/x_{f}$ is an increasing function of random variables $Z_{\Gamma(v)\setminus L}$. These sets $\Gamma(v)\setminus L$ are disjoint since the right-endpoints of edges in $L$ are all distinct. Since random variables $Z_{L^{\prime}}$ are NA, Theorem 4 yields
$\Pr\Bigl(\bigwedge_{(u,v)\in L}Z_{(u,v)}/x_{u,v}=\min_{f\in\Gamma(v)}Z_{f}/x_{f}\Bigr)\leq\prod_{(u,v)\in L}\Pr\bigl(Z_{(u,v)}/x_{(u,v)}=\min_{f\in\Gamma(v)}Z_{f}/x_{f}\bigr)$
For any edge $e=(u,v)\in L$, the variables $Z_{f}:f\in\Gamma(v)\setminus\{e\}$ are independent unit-rate Exponentials. By standard facts about Exponential random variables, this implies that $Z^{\prime}:=\min_{f\in\Gamma(v)\setminus\{e\}}Z_{f}/x_{f}$ is an Exponential random variable with rate $\sum_{f\in\Gamma(v)\setminus\{e\}}x_{f}=1-x_{e}$. The probability that $Z^{\prime}>Z_{e}/x_{e}$ is precisely $\mathrm{e}^{-(1-x_{e})\cdot Z_{e}/x_{e}}=\mathrm{e}^{(1-1/x_{e})Z_{e}}$. ∎
Proposition 12.
Property (A3) holds.
Proof.
Consider a stable edge set $L\subseteq E$. By iterated expectations with respect to random variable $Z_{L}$ and Lemma 11, we have
${\mathbb{E}}\Bigl[\prod_{e\in L}X_{e}\Bigr]={\mathbb{E}}_{Z_{L}}\Bigl[{\mathbb{E}}\bigl[\prod_{e\in L}X_{e}\mid Z_{L}\bigr]\Bigr]\leq{\mathbb{E}}\Bigl[\prod_{e\in L}\mathrm{e}^{(1-1/x_{e})Z_{e}}\Bigr]$
Each term $\mathrm{e}^{(1-1/x_{e})Z_{e}}$ in this product is a decreasing function of random variable $Z_{e}$. Since the variables $Z_{L}$ are NA, Theorem 4 gives:
${\mathbb{E}}\Bigl[\prod_{e\in L}\mathrm{e}^{(1-1/x_{e})Z_{e}}\Bigr]\leq\prod_{e\in L}{\mathbb{E}}\bigl[\mathrm{e}^{(1-1/x_{e})Z_{e}}\bigr]$
Here, for an edge $e$, we have ${\mathbb{E}}[\mathrm{e}^{(1-1/x_{e})Z_{e}}]=\int_{z=0}^{\infty}{\mathrm{e}}^{-z}\cdot\mathrm{e}^{(1-1/x_{e})z}\ \mathrm{d}z=x_{e}.$
Theorem 13.
For any edges $e_{1}=(u,v_{1}),e_{2}=(u,v_{2})$ with the same left-node $u$, there holds ${\mathbb{E}}[X_{e_{1}}X_{e_{2}}]\leq x_{e_{1}}x_{e_{2}}\cdot(1-\Phi(x_{e_{1}},x_{e_{2}};\rho_{e_{1}},\rho_{e_{2}}))$.
Proof.
Let us write $X_{i}=X_{e_{i}},x_{i}=x_{e_{i}},\rho_{i}=\rho_{e_{i}},Z_{i}=Z_{e_{i}}$ for $i=1,2$. By Lemma 11 applied to stable-set $L=\{e_{1},e_{2}\}$ we have:
${\mathbb{E}}[X_{1}X_{2}\mid Z_{1},Z_{2}]$ $\leq\mathrm{e}^{(1-1/x_{1})Z_{1}+(1-1/x_{2})Z_{2}}$
Observe that values $Z_{1},Z_{2}$ are simultaneously generated by $\textsc{CorrelatedExponential}(\vec{\rho}^{(u)})$. By applying Lemma 8 with $q_{1}=1-1/x_{1},q_{2}=1-1/x_{2}$, we have
${\mathbb{E}}[\mathrm{e}^{(1-1/x_{1})Z_{1}+(1-1/x_{2})Z_{2}}]=x_{1}x_{2}\cdot\Bigl(1-\frac{((1-\rho_{1})^{1-1/x_{1}}-1)((1-\rho_{2})^{1-1/x_{2}}-1)}{(1-\rho_{1})^{1-1/x_{1}}(1-\rho_{2})^{1-1/x_{2}}+\rho_{1}+\rho_{2}-1}\Bigr).\qed$
As we have mentioned, there is no single optimal choice for the vector $\vec{\rho}$. One attractive option is to simply set $\vec{\rho}$ proportional to $\vec{p}$, which leads to the following simple version of property (A4):
Lemma 14.
For $t\geq 4/3$ and $x_{1},x_{2}\in(0,1)$, there holds
$\Phi(x_{1},x_{2};tx_{1},tx_{2})\geq 1-\frac{\mathrm{e}^{t}-1}{\mathrm{e}^{t}+1}+0.57(x_{1}+x_{2})\max\{0,1/t-0.45\}$
The proof of Lemma 14 involves significant numerical analysis; we defer it to Appendix A. In the scheduling algorithm, most edges will set $\rho_{e}\propto x_{e}$, and use the bound in Lemma 14. However, a few edges will use a different value of $\rho_{e}$. This is a good illustration of the flexibility of the bipartite dependent-rounding scheme.
Another nice choice, which can work for more general graphs, is the following:
Proposition 15.
For $t>0$ and $x_{1},x_{2}\in(0,1)$, there holds
$\Phi(x_{1},x_{2};1-\mathrm{e}^{-x_{1}t},1-\mathrm{e}^{-x_{2}t})=1-\frac{(\mathrm{e}^{t}-1)(\mathrm{e}^{x_{1}t}+\mathrm{e}^{x_{2}t})}{\mathrm{e}^{2t}-\mathrm{e}^{x_{1}t}-\mathrm{e}^{x_{2}t}+\mathrm{e}^{(x_{1}+x_{2})t}}\geq 1-\frac{\mathrm{e}^{x_{1}t}+\mathrm{e}^{x_{2}t}}{1+\mathrm{e}^{t}}$
Proof.
Let $\rho_{i}=1-\mathrm{e}^{x_{i}t}$ for $i=1,2$, and let $\eta=\mathrm{e}^{t}$. With some algebraic simplifications we get
$\Phi(x_{1},x_{2};\rho_{1},\rho_{2})=1-\frac{(\mathrm{e}^{t}-1)(\mathrm{e}^{x_{1}t}+\mathrm{e}^{x_{2}t})}{\mathrm{e}^{2t}-\mathrm{e}^{x_{1}t}-\mathrm{e}^{x_{2}t}+\mathrm{e}^{(x_{1}+x_{2})t}}=1-\frac{(\eta-1)(\eta^{x_{1}}+\eta^{x_{2}})}{\eta^{2}-\eta^{x_{1}}-\eta^{x_{2}}+\eta^{x_{1}+x_{2}}}$
Furthermore, we can observe that
$\frac{(\eta-1)(\eta^{x_{1}}+\eta^{x_{2}})}{\eta^{2}-\eta^{x_{1}}-\eta^{x_{2}}+\eta^{x_{1}+x_{2}}}=\frac{(\eta-1)(\eta^{x_{1}}+\eta^{x_{2}})}{(\eta^{x_{1}}-1)(\eta^{x_{2}}-1)+(\eta^{2}-1)}\leq\frac{(\eta-1)(\eta^{x_{1}}+\eta^{x_{2}})}{(\eta^{2}-1)}=\frac{\eta^{x_{1}}+\eta^{x_{2}}}{\eta+1}.\qed$
4 Scheduling to minimize completion time
Our general algorithm can be summarized as follows:
  • Solve the SDP relaxation, obtaining estimates $x^{(i)}_{j,j^{\prime}}$ for each machine $i$ and pair of jobs $j,j^{\prime}$. Roughly speaking, $x^{(i)}_{j,j^{\prime}}$ represents the fractional extent to which jobs $j,j^{\prime}$ are simultaneously scheduled on machine $i$. The case with $j=j^{\prime}$ plays an especially important role, in which case we write simply $x^{(i)}_{j}$.
  • Based on the fractional solution $\vec{x}$, partition the jobs on each machine $i$ into clusters $\mathcal{C}^{(i)}_{1},\dots,\mathcal{C}^{(i)}_{t}$.
  • Run DepRound to obtain rounded variables $X^{(i)}_{j}\in\{0,1\}$, with $\sum_{i}X^{(i)}_{j}=1$ for all jobs $j$. Job $j$ is assigned to the machine $i$ with $X^{(i)}_{j}=1$.
  • Schedule jobs assigned to each machine $i$ in non-increasing order of Smith ratio $\sigma^{(i)}_{j}=w_{j}/p^{(i)}_{j}$.
We will not modify the first or last steps in any way; they are discussed in more detail in Section 4.1 next. The difference is how we implement the second step and third steps (partitioning and rounding the jobs). Specifically, given a fractional solution $x$, we form our clusters as follows:
1Define parameters $\pi=3.9,\theta=0.555,\tau=0.604$.
2Draw a random variable $P_{\text{offset}}$ uniformly at random from $[0,1]$.
3for each machine $i$ do
4
5  Partition the jobs into processing time classes $\mathcal{P}^{(i)}_{k}=\{j:P_{\text{offset}}+\frac{\log p^{(i)}_{j}}{\log\pi}\in[k,k+1)\}$.
6 for each class $\mathcal{P}^{(i)}_{k}$ do
7
8     Initialize cluster index $\ell=1$ and set $\mathcal{C}^{(i)}_{k,1}$$\emptyset$.
9     Sort the jobs in $\mathcal{P}^{(i)}_{k}$ with $x^{(i)}_{j}>0$ in non-increasing order of Smith ratio as $j_{1},j_{2},\dots,j_{s}$.
10    for $t=1,\dots,s$ and each job $j=j_{t}$ do
11
12        Set $\tilde{\rho}^{(i)}_{j}=\min\{x^{(i)}_{j},\tau-\sum_{j^{\prime}\in\mathcal{C}^{(i)}_{k,\ell}}x^{(i)}_{j^{\prime}}\}$
13        Update $\mathcal{C}^{(i)}_{k,\ell}\leftarrow\mathcal{C}^{(i)}_{k,\ell}\cup\{j_{t}\}$.
14       if $\sum_{j\in\mathcal{C}^{(i)}_{k,\ell}}x^{(i)}_{j}\geq\theta$ then
15
16           Update $\ell\leftarrow\ell+1$ and initialize the new cluster $\mathcal{C}^{(i)}_{k,\ell}$$\emptyset$
17
18
19     For each cluster $k,\ell$ and every job $j\in\mathcal{C}^{(i)}_{k,\ell}$ set $\rho^{(i)}_{j}=\frac{\tilde{\rho}^{(i)}_{j}}{\sum_{j^{\prime}\in\mathcal{C}^{(i)}_{k,\ell}}\tilde{\rho}^{(i)}_{j^{\prime}}}$.
20
21  Run $\textsc{DepRound}(\rho,x)$ to convert the fractional solution $x$ into an integral solution $X$.
22
Algorithm 3 Clustering the jobs
So each machine has a single “open” cluster for each processing-time quantization class at a time. If, after adding the job to the open cluster, the cluster size becomes at least $\theta$, then we close it and open a new one. The correlation parameter $\rho^{(i)}_{j}$ is usually chosen to be proportional to $x^{(i)}_{j}$; the one exception is that the final job that closes out a cluster may need to choose a smaller parameter $\rho^{(i)}_{j}$. We say a job $j$ is truncated on a machine $i$ if $\tilde{\rho}_{j}^{(i)}<x^{(i)}_{j}$; note that, in this case, every other job $j^{\prime}$ in the cluster has $\rho_{j^{\prime}}^{(i)}=x^{(i)}_{j}/\tau$.
[…]
Proposition 16.
Algorithm DepRound can be used to round the fractional solution $x$ to an integral solution $X$, where for any machine $i$ and distinct jobs $j,j^{\prime}$, it satisfies
${\mathbb{E}}[X^{(i)}_{j}]=x^{(i)}_{j},\qquad{\mathbb{E}}[X^{(i)}_{j}X^{(i)}_{j^{\prime}}]\leq x^{(i)}_{j}x^{(i)}_{j^{\prime}}$
and for any machine $i$ and any cluster $k,\ell$ and distinct jobs $j,j^{\prime}\in\mathcal{C}^{(i)}_{k,\ell}$, it has
${\mathbb{E}}[X^{(i)}_{j}X^{(i)}_{j^{\prime}}]\leq(1-\phi^{(i)}_{j,j^{\prime}})\cdot x^{(i)}_{j}x^{(i)}_{j^{\prime}}\quad\text{ for $\phi^{(i)}_{j,j^{\prime}}:=\Phi(x^{(i)}_{j},x^{(i)}_{j^{\prime}};\rho^{(i)}_{j},\rho^{(i)}_{j^{\prime}})$}$
Proof.
Apply Theorem 1 to the graph with left-nodes $\mathcal{C}^{(i)}_{k,\ell}$ and right-nodes $\mathcal{J}$; for each machine $i$ and job $j$ where $j\in\mathcal{C}^{(i)}_{k,\ell}$, it has an edge $e=(\mathcal{C}^{(i)}_{k,\ell},j)$ with $x_{e}=x^{(i)}_{j},\rho_{e}=\rho^{(i)}_{j},X^{(i)}_{j}=X_{e}$. Note that $\sum_{e\in\Gamma(j)}x_{e}=\sum_{e\in\mathcal{C}^{(i)}_{k,\ell}}\rho_{e}=1$.
By Property (A1), every job is assigned to exactly one machine. By Property (A2), we have ${\mathbb{E}}[X^{(i)}_{j}]=x^{(i)}_{j}$. For any pair of jobs $j,j^{\prime}$ on a machine $i$, note that the corresponding edge-set $\{(\mathcal{C}^{(i)}_{k,\ell},j),(\mathcal{C}^{(i)}_{k^{\prime},\ell^{\prime}},j^{\prime})\}$ is stable. So Property (A3) gives ${\mathbb{E}}[X^{(i)}_{j}X^{(i)}_{j^{\prime}}]\leq x^{(i)}_{j}x^{(i)}_{j^{\prime}}$. Also, by Property (A4), every pair of jobs $j,j^{\prime}$ in a cluster $\mathcal{C}^{(i)}_{k,\ell}$ has ${\mathbb{E}}[X^{(i)}_{j}X^{(i)}_{j^{\prime}}]\leq(1-\phi^{(i)}_{j,j^{\prime}})x^{(i)}_{j}x^{(i)}_{j^{\prime}}$. ∎
For each class $\mathcal{P}^{(i)}_{k}$ and each job $j\in\mathcal{P}^{(i)}_{k}$, we define
$P_{k}=\pi^{k-P_{\text{offset}}},\qquad H^{(i)}_{j}=p^{(i)}_{j}/P_{k}.$
Note that $H^{(i)}_{j}\in[1,\pi]$ for all $j$. The key property we exploit is that each random value $\log_{\pi}H^{(i)}_{j}$ is uniformly distributed in $[0,1]$. This leads to the following formula:
Observation 17.
For any job $j$ and any function $\Psi:[1,\pi]\rightarrow\mathbb{R}$, we have
${\mathbb{E}}[\Psi(H_{j})]=\frac{1}{\log\pi}\int_{h=1}^{\pi}\frac{\Psi(h)}{h}\ \mathrm{d}h$
4.1 The SDP relaxation
Here, we provide a brief summary of the SDP relaxation and its properties. See [2] for more details.
For a single machine, there is a simple heuristic to minimize weighted completion time: namely, jobs should be ordered in non-increasing order of their Smith ratio $\sigma^{(i)}_{j}$. To simplify notation, let us suppose that all the ratios $\sigma^{(i)}_{j}$ are distinct. (This can be achieved without loss of generality by adding infinitesimal noise to each $p^{(i)}_{j}$.) Thus, the overall weighted completion time would be
$\sum_{\begin{subarray}{c}j,j^{\prime}\text{assigned to machine $i$}\\ \sigma^{(i)}_{j^{\prime}}\leq\sigma^{(i)}_{j}\end{subarray}}w_{j}p^{(i)}_{j^{\prime}}$
For each machine $i$, let us further define a symmetric $(|\mathcal{J}|+1)\times(|\mathcal{J}|+1)$ matrix $\mathbf{X}^{(i)}$ as follows: we set $\mathbf{X}^{(i)}_{0,0}=1$, we set $\mathbf{X}^{(i)}_{0,j}=\mathbf{X}^{(i)}_{j,0}=x^{(i)}_{j}$, and we set $\mathbf{X}^{(i)}_{j,j^{\prime}}=x^{(i)}_{j,j^{\prime}}$ for all pairs of jobs $j,j^{\prime}$. This motivates the following semidefinite-programming (SDP) relaxation:
maximize $\qquad\qquad\sum_{i\in\mathcal{M}}\sum_{j\in\mathcal{J}}w_{j}\sum_{\begin{subarray}{c}j^{\prime}\in\mathcal{J}:\sigma^{(i)}_{j^{\prime}}\leq\sigma^{(i)}_{j}\end{subarray}}p^{(i)}_{j^{\prime}}x^{(i)}_{j,j^{\prime}}$
[…]
$\qquad\qquad\sum_{i\in\mathcal{M}}x^{(i)}_{j}=1\qquad\text{for all $j$}$
Each matrix $\mathbf{X}^{(i)}$ is symmetric and positive-semidefinite
As described in [2], this relaxation can be solved in polynomial time. Furthermore, given an integral solution $X^{(i)}_{j}\in\{0,1\}^{\mathcal{J}}$, there is a corresponding SDP solution defined by $x^{(i)}_{j,j^{\prime}}=X^{(i)}_{j}X^{(i)}_{j^{\prime}}$. Our goal is to convert the fractional assignments $x^{(i)}_{j}$ into integral assignments $X^{(i)}_{j}$. We quote the following key results of [2] concerning the semidefinite program.
Theorem 18 ([2]).
For any machine $i^{*}$, suppose the jobs are sorted $1,\dots,n$ in non-increasing order of Smith ratio $\sigma^{(i^{*})}_{j}$ along with a dummy job $n+1$ with $w_{n+1}=\sigma^{(i^{*})}_{n+1}=0$. Then the completion time on machine $i^{*}$ is
$\sum_{j^{*}=1}^{n}(\sigma^{(i^{*})}_{j^{*}}-\sigma^{(i^{*})}_{j^{*}+1})Z^{(i^{*},j^{*})}$
while the contribution to the SDP objective function corresponding to machine $i^{*}$ is given by
$\sum_{j^{*}=1}^{n}(\sigma^{(i^{*})}_{j^{*}}-\sigma^{(i^{*})}_{j^{*}+1})\operatorname{LB}^{(i^{*},j^{*})}$
where for each job $j^{*}$ we define
$Z^{(i^{*},j^{*})}$ $=\frac{1}{2}\Bigl(\sum_{j=1}^{j^{*}}X_{j}^{(i^{*})}(p_{j}^{(i^{*})})^{2}+\sum_{j=1}^{j^{*}}\sum_{j^{\prime}=1}^{j^{*}}X_{j}^{(i^{*})}X_{j^{\prime}}^{(i^{*})}p_{j}^{(i^{*})}p_{j^{\prime}}^{(i^{*})}\Bigr)$
$\operatorname{LB}^{(i^{*},j^{*})}$ $=\frac{1}{2}\Bigl(\sum_{j=1}^{j^{*}}x_{j}^{(i^{*})}(p_{j}^{(i^{*})})^{2}+\sum_{j=1}^{j^{*}}\sum_{j^{\prime}=1}^{j^{*}}x^{(i^{*})}_{j,j^{\prime}}p^{(i^{*})}_{j}p^{(i^{*})}_{j^{\prime}}\Bigr)$
Corollary 19.
If ${\mathbb{E}}[Z^{(i^{*},j^{*})}]\leq\eta\cdot\operatorname{LB}^{(i^{*},j^{*})}$ for all machines $i^{*}$ and jobs $j^{*}$, then the resulting schedule is an $\eta$-approximation in expectation, and the SDP has integrality gap at most $\eta$.
4.2 Focusing on a single machine and job.
The main consequence of Corollary 19 is that we can focus on a single machine $i^{*}$ and single job $j^{*}$ and ignore all the job weights. For the remainder of the analysis, we suppose $i^{*},j^{*}$ are fixed, and we omit all superscripts $(i^{*},j^{*})$, for example we write $X_{j}$ instead of $X^{(i^{*})}_{j}$.
We define $\mathcal{J}^{*}$ to be the set of jobs $j$ with $x_{j}>0$ and $\sigma_{j}\geq\sigma_{j^{*}}$. As a point of notation, any sum of the form $\sum_{j}$ should be taken to range over $j\in\mathcal{J}^{*}$ unless stated otherwise. Likewise, in a sum of the form $\sum_{j,j^{\prime}}$, we view $j,j^{\prime}$ as an ordered pair of jobs in $\mathcal{J}^{*}$; there are separate summands for $j,j^{\prime}$ and for $j^{\prime},j$ and we also allow $j=j^{\prime}$.
For each class $\mathcal{P}_{k}$ we define $\mathcal{P}^{*}_{k}=\mathcal{P}_{k}\cap\mathcal{J}^{*}$ and for each cluster $\mathcal{C}_{k,\ell}$, we define $\mathcal{C}^{*}_{k,\ell}=\mathcal{C}_{k,\ell}\cap\mathcal{J}^{*}$. Within each class $\mathcal{P}_{k}$, the final opened cluster $\mathcal{C}^{*}_{k,\ell}$ after processing job $j^{*}$ is called the leftover cluster for $k$, and denoted by $\mathcal{C}^{*}_{k,\text{left}}$.
With these conventions, we can write $Z=Z^{(i^{*},j^{*})}$ and $\operatorname{LB}=\operatorname{LB}^{(i^{*},j^{*})}$ more compactly as:
$Z$ $=\frac{1}{2}\Bigl(\sum_{j}X_{j}p_{j}^{2}+\sum_{j,j^{\prime}}X_{j}X_{j^{\prime}}p_{j}p_{j^{\prime}}\Bigr)=\sum_{j}X_{j}p_{j}^{2}+\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}:j^{\prime}\neq j\end{subarray}}X_{j}X_{j^{\prime}}p_{j}p_{j^{\prime}}$
$\operatorname{LB}$ $=\frac{1}{2}\Bigl(\sum_{j}x_{j}p_{j}^{2}+\sum_{j,j^{\prime}}x_{j,j^{\prime}}p_{j}p_{j^{\prime}}\Bigr)\quad=\sum_{j}x_{j}p_{j}^{2}+\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}:j^{\prime}\neq j\end{subarray}}x_{j,j^{\prime}}p_{j}p_{j^{\prime}}$
Following [2], we further define two important quantities for measuring the approximation ratio.
$L=\sum_{j\in\mathcal{J}^{*}}x_{j}p_{j},\qquad Q=\sum_{j\in\mathcal{J}^{*}}x_{j}p_{j}^{2}$
Theorem 20.
For any vector $y\in[0,1]^{\mathcal{J}^{*}}$ there holds [^4] Theorem 20 was shown in [2] for an integral vector $y\in\{0,1\}^{\mathcal{J}^{*}}$, but we will need the slightly generalized result for our analysis.
$\operatorname{LB}\geq\frac{1}{2}\Bigl(Q+\sum_{j}y_{j}x_{j}p_{j}^{2}+\Bigl(L-\sum_{j}(1-\sqrt{1-y_{j}})\cdot x_{j}p_{j}\Bigr)^{2}\Bigr)$
Proof.
Define $z_{j}=\sqrt{1-y_{j}}$. We decompose the expression for $\operatorname{LB}$ as:
$\frac{1}{2}\Bigl(\sum_{j}x_{j}p_{j}^{2}+\sum_{j,j^{\prime}}x_{j,j^{\prime}}p_{j}p_{j^{\prime}}\Bigr)=\frac{1}{2}\Bigl(\sum_{j}x_{j}p_{j}^{2}+\sum_{j,j^{\prime}}(1-z_{j}z_{j^{\prime}})x_{j,j^{\prime}}p_{j}p_{j^{\prime}}+\sum_{j,j^{\prime}}x_{j,j^{\prime}}z_{j}z_{j^{\prime}}p_{j}p_{j^{\prime}}\Bigr)$ (4)
The sum $\sum_{j,j^{\prime}}(1-z_{j}z_{j^{\prime}})x_{j,j^{\prime}}p_{j}p_{j^{\prime}}$ can be lower-bounded by only including the terms with $j=j^{\prime}$, which contribute $\sum_{j}(1-z_{j}^{2})x_{j}p_{j}^{2}=\sum_{j}y_{j}x_{j}p_{j}^{2}$.
For the sum $\sum_{j,j^{\prime}}x_{j,j^{\prime}}z_{j}z_{j^{\prime}}p_{j}p_{j^{\prime}}$, let $\mu=\sum_{j}z_{j}x_{j}p_{j}=L-\sum_{j}(1-\sqrt{1-y_{j}})x_{j}p_{j}$. Define a vector $v\in\mathbb{R}^{|\mathcal{J}|+1}$ by setting $v_{0}=-\mu,v_{j}=z_{j}p_{j}$ for $j\in\mathcal{J}^{*}$, and $v_{j}=0$ for $j\in\mathcal{J}\setminus\mathcal{J}^{*}$. The SDP relaxation ensures that $v^{\top}\mathbf{X}v\geq 0$. We calculate:
$v^{\top}\mathbf{X}v$ $=\mathbf{X}_{0,0}v_{0}^{2}+2\sum_{j}\mathbf{X}_{0,j}v_{j}v_{0}+\sum_{j,j^{\prime}}\mathbf{X}_{j,j^{\prime}}v_{j}v_{j^{\prime}}$
$=\mu^{2}-2\sum_{j}x_{j}z_{j}p_{j}\mu+\sum_{j,j^{\prime}}x_{j,j^{\prime}}z_{j}z_{j^{\prime}}p_{j}p_{j^{\prime}}=-\mu^{2}+\sum_{j,j^{\prime}}x_{j,j^{\prime}}z_{j}z_{j^{\prime}}p_{j}p_{j^{\prime}}$
So $\sum_{j,j^{\prime}}x_{j,j^{\prime}}z_{j}z_{j^{\prime}}p_{j}p_{j}^{\prime}\geq\mu^{2}$. Substituting into Eq. (4) gives:
$\operatorname{LB}\geq\frac{1}{2}\Bigl(Q+\sum_{j}y_{j}x_{j}p_{j}^{2}+\Bigl(L-\sum_{j}(1-\sqrt{1-y_{j}})x_{j}p_{j}\Bigr)^{2}\Bigr).\qed$
Corollary 21 ([2]).
There holds $\operatorname{LB}\geq\max\{Q,\tfrac{1}{2}(Q+L^{2})\}$.
Proof.
Apply Theorem 20 with respectively vectors $y=\vec{1}$ and $y=\vec{0}$. ∎
At this point, we can give a simple explanation of how the algorithm achieves a $1.5$-approximation where we only assume nonpositive correlation among the variable $X_{j}$. To get a better approximation ratio we need to show there is strong negative correlation among them.
Proposition 22 (assuming nonpositive correlation only).
There holds ${\mathbb{E}}[Z]\leq 3/2\cdot\operatorname{LB}$.
Proof.
For any jobs $j,j^{\prime}$ we have ${\mathbb{E}}[X_{j}]=x_{j}$ and ${\mathbb{E}}[X_{j}X_{j^{\prime}}]\leq x_{j}x_{j^{\prime}}$. So
${\mathbb{E}}[Z]\leq$ $\sum_{j}p_{j}^{2}x_{j}+\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}:j^{\prime}\neq j\end{subarray}}p_{j}p_{j^{\prime}}x_{j}x_{j^{\prime}}=\sum_{j}p_{j}^{2}(x_{j}-x_{j}^{2}/2)+\frac{1}{2}\sum_{j^{\prime},j}p_{j}p_{j^{\prime}}x_{j}x_{j^{\prime}}\leq Q+\frac{1}{2}L^{2}$
On the other hand, Corollary 21 gives
$\operatorname{LB}\geq\max\{Q,\tfrac{1}{2}(Q+L^{2})\}\geq\frac{1}{3}Q+\frac{2}{3}(\tfrac{1}{2}(Q+L^{2}))=\frac{2}{3}Q+\frac{1}{3}L^{2}.\qed$
5 Determining the approximation ratio
We now begin the arduous computation of the approximation ratio. The analysis has three main steps. First, we compute an upper bound on the expected value of the solution returned by the rounding algorithm. Second, we compute a lower bound on the objective function of the relaxation. Finally, we combine these two estimates.
We will introduce parameters $\kappa,\beta,\delta,\gamma$ to bound various internal functions, along with related numerical constants $c_{0},\dots,c_{6}$. All calculations are carried out using exact arithmetic in the Mathematica computer algebra system; some specific calculation details are deferred to Appendix B. We emphasize that these parameters are not used directly in the algorithm itself.
5.1 The algorithm upper-bound
As a starting point, we have the following identity:
Lemma 23.
If we condition on random variable $P_{\text{offset}}$, then the expectation over the procedure DepRound satisfies
${\mathbb{E}}[Z\mid P_{\text{offset}}]\leq Q+\frac{L^{2}}{2}-\sum_{k,\ell}P_{k}^{2}B_{k,\ell}$
where for each cluster $k,\ell$ we define the “bonus term”
$B_{k,\ell}=\frac{1}{2}\Bigl(\sum_{j\in\mathcal{C}^{*}_{k,\ell}}x_{j}^{2}H_{j}^{2}+\sum_{\begin{subarray}{c}j,j^{\prime}\in\mathcal{C}^{*}_{k,\ell}:j\neq j^{\prime}\end{subarray}}\phi_{j,j^{\prime}}x_{j}x_{j^{\prime}}H_{j}H_{j^{\prime}}\Bigr)$
where recall that $\phi_{j,j^{\prime}}=\Phi(x_{j},x_{j^{\prime}};\rho_{j},\rho_{j^{\prime}}).$
Proof.
All calculations in this proof are conditioned on $P_{\text{offset}}$. We have:
${\mathbb{E}}[Z]={\mathbb{E}}\Bigl[\sum_{j}p_{j}^{2}X_{j}+\frac{1}{2}\sum_{j^{\prime}\neq j}p_{j}p_{j^{\prime}}X_{j}X_{j^{\prime}}\Bigr]$
Here ${\mathbb{E}}[X_{j}]=x_{j}$, and for any pair $j,j^{\prime}$, we have ${\mathbb{E}}[X_{j}X_{j^{\prime}}]\leq x_{j}x_{j^{\prime}}$. Also, for any cluster $k,\ell$ and distinct jobs $j,j^{\prime}\in\mathcal{C}^{*}_{k,\ell}$, we have ${\mathbb{E}}[X_{j}X_{j^{\prime}}]\leq(1-\phi_{j,j^{\prime}})x_{j}x_{j^{\prime}}$ by Proposition 16. So:
${\mathbb{E}}[Z]\leq\sum_{j}p_{j}^{2}x_{j}+\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}:j\neq j^{\prime}\end{subarray}}p_{j}p_{j^{\prime}}x_{j}x_{j^{\prime}}-\frac{1}{2}\sum_{k,\ell}\sum_{\begin{subarray}{c}j,j^{\prime}\in\mathcal{C}^{*}_{k,\ell}:j\neq j^{\prime}\end{subarray}}\phi_{j,j^{\prime}}x_{j}x_{j^{\prime}}p_{j}p_{j^{\prime}}$
$=\sum_{j}p_{j}^{2}(x_{j}-x_{j}^{2}/2)+\frac{1}{2}\sum_{j^{\prime},j}p_{j}p_{j^{\prime}}x_{j}x_{j^{\prime}}-\frac{1}{2}\sum_{k,\ell}\sum_{\begin{subarray}{c}j,j^{\prime}\in\mathcal{C}^{*}_{k,\ell}:j\neq j^{\prime}\end{subarray}}\phi_{j,j^{\prime}}x_{j}x_{j^{\prime}}p_{j}p_{j^{\prime}}$
$=Q+\frac{L^{2}}{2}-\frac{1}{2}\sum_{k,\ell}\Bigl(\sum_{j\in\mathcal{C}^{*}_{k,\ell}}p_{j}^{2}x_{j}^{2}+\negthickspace\negthickspace\sum_{\begin{subarray}{c}j,j^{\prime}\in\mathcal{C}^{*}_{k,\ell}:j\neq j^{\prime}\end{subarray}}\negthickspace\negthickspace\phi_{j,j^{\prime}}x_{j}x_{j^{\prime}}p_{j}p_{j^{\prime}}\Bigr)$
It remains to observe that, for any cluster $k,\ell$ and job $j\in\mathcal{C}^{*}_{k,\ell}$, we have $p_{j}=H_{j}P_{k}$, and so
$\sum_{j\in\mathcal{C}^{*}_{k,\ell}}p_{j}^{2}x_{j}^{2}+\negthickspace\negthickspace\sum_{\begin{subarray}{c}j,j^{\prime}\in\mathcal{C}^{*}_{k,\ell}:j\neq j^{\prime}\end{subarray}}\negthickspace\negthickspace\phi_{j,j^{\prime}}x_{j}x_{j^{\prime}}p_{j}p_{j^{\prime}}=\sum_{j\in\mathcal{C}^{*}_{k,\ell}}(P_{k}^{2}H_{j}^{2})x_{j}^{2}+\negthickspace\negthickspace\sum_{\begin{subarray}{c}j,j^{\prime}\in\mathcal{C}^{*}_{k,\ell}:j\neq j^{\prime}\end{subarray}}\negthickspace\negthickspace\phi_{j,j^{\prime}}x_{j}x_{j^{\prime}}(P_{k}H_{j})(P_{k}H_{j^{\prime}})$
$=P_{k}^{2}\Bigl(\sum_{j\in\mathcal{C}^{*}_{k,\ell}}x_{j}^{2}H_{j}^{2}+\negthickspace\negthickspace\sum_{\begin{subarray}{c}j,j^{\prime}\in\mathcal{C}^{*}_{k,\ell}:j\neq j^{\prime}\end{subarray}}\negthickspace\negthickspace\phi_{j,j^{\prime}}x_{j}x_{j^{\prime}}H_{j}H_{j^{\prime}}\Bigr)=2P_{k}^{2}B_{k,\ell}.\qed$
The main focus of the analysis is to show that $B_{k,\ell}$ is large on average. For intuition, keep in mind the extremal case when all untruncated jobs have infinitesimal mass (i.e. $x_{j}^{2}\approx 0$). The truncated jobs require some special analysis. We also need to show that non-infinitesimal mass can only increase the value $B_{k,\ell}$.
Lemma 24.
For a cluster $k,\ell$, let $\mathcal{U},\mathcal{T}$ be the set of untruncated jobs and truncated jobs respectively in $\mathcal{C}^{*}_{k,\ell}$ and define parameters as follows:
$\lambda=\sum_{j\in\mathcal{C}^{*}_{k,\ell}}\tilde{\rho}_{j},\quad a=\frac{\mathrm{e}^{1/\lambda}-1}{\mathrm{e}^{1/\lambda}+1},\quad r=\sum_{j\in\mathcal{U}}x_{j},\quad s=\sum_{j\in\mathcal{U}}x_{j}H_{j},\quad y=\sum_{j\in\mathcal{T}}x_{j},\quad d=\sum_{j\in\mathcal{T}}x_{j}H_{j}$
We have:
$B_{k,\ell}\geq d^{2}/2+s/2\cdot\inf_{x\in(0,r]}\Bigl(x+(a-c_{0}x)(s-x)+2d\Phi(x,y;x/\tau,1-r/\tau)\Bigr)$
for constant $c_{0}:=0.17556$.
Proof.
We first claim that there holds:
$\sum_{\begin{subarray}{c}j,j^{\prime}\in\mathcal{U}:j\neq j^{\prime}\end{subarray}}\phi_{j,j^{\prime}}x_{j}x_{j^{\prime}}H_{j}H_{j^{\prime}}\geq\sum_{j\in\mathcal{U}}x_{j}H_{j}(a-c_{0}x_{j})(s-x_{j})\quad$ (5)
For, each untruncated job $j$ has $\rho_{j}=\tilde{\rho}_{j}/\lambda=x_{j}/\lambda$ where $\lambda\leq\tau\leq 3/4$. Because we only add jobs with non-zero mass to clusters, it also must have $x_{j}>0$. By Lemma 14, any pair $j,j^{\prime}\in\mathcal{U}$ thus has $\phi_{j,j^{\prime}}\geq a-b(x_{j}+x_{j^{\prime}})$ for $b=0.57\max\{0,\lambda-0.45\}\leq 0.57(\tau-0.45)=c_{0}/2$. So:
$\sum_{\begin{subarray}{c}j,j^{\prime}\in\mathcal{U}:j\neq j^{\prime}\end{subarray}}\phi_{j,j^{\prime}}x_{j}x_{j^{\prime}}H_{j}H_{j^{\prime}}\geq\sum_{\begin{subarray}{c}j,j^{\prime}\in\mathcal{U}:j\neq j^{\prime}\end{subarray}}\bigl(a-c_{0}/2\cdot(x_{j}+x_{j^{\prime}})\bigr)x_{j}x_{j^{\prime}}H_{j}H_{j^{\prime}}$
$\qquad=\sum_{j,j^{\prime}\in\mathcal{U}:j\neq j^{\prime}}\bigl(a-c_{0}x_{j}\bigr)x_{j}x_{j^{\prime}}H_{j}H_{j^{\prime}}\qquad\text{(by symmetry)}$
$\qquad=\sum_{j\in\mathcal{U}}x_{j}H_{j}\bigl(a-c_{0}x_{j}\bigr)\bigl(-x_{j}H_{j}+\sum_{j^{\prime}\in\mathcal{U}}x_{j^{\prime}}H_{j^{\prime}}\bigr)=\sum_{j\in\mathcal{U}}x_{j}H_{j}(a-c_{0}x_{j})(-x_{j}H_{j}+s)$
Finally, since $a\geq\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}\geq 0.68$ and $H_{j}\geq 1$, we have $(a-c_{0}x_{j})(s-x_{j}H_{j})\geq(a-c_{0}x_{j})(s-x_{j})$, which establishes the bound (5).
Now, noting that $|\mathcal{T}|\leq 1$, we can decompose $B_{k,\ell}$ as follows:
$2B_{k,\ell}$ $=\sum_{t\in\mathcal{T}}x_{t}^{2}H_{t}^{2}+\sum_{j\in\mathcal{U}}x_{j}^{2}H_{j}^{2}+\sum_{\begin{subarray}{c}j,j^{\prime}\in\mathcal{U}:j\neq j^{\prime}\end{subarray}}\phi_{j,j^{\prime}}x_{j}x_{j^{\prime}}H_{j}H_{j^{\prime}}+2\sum_{\begin{subarray}{c}j\in\mathcal{U},t\in\mathcal{T}\end{subarray}}\phi_{j,t}x_{j}x_{t}H_{j}H_{t}$
$\geq\sum_{t\in\mathcal{T}}x_{t}^{2}H_{t}^{2}+\sum_{j\in\mathcal{U}}x_{j}H_{j}\Bigl(x_{j}+(a-c_{0}x_{j})(s-x_{j})+2\sum_{t\in\mathcal{T}}\phi_{j,t}x_{t}H_{t}\Bigr)$
If $\mathcal{C}^{*}_{k,\ell}$ has a truncated job $t$ with $\rho_{t}=1-r/\tau$, then for any job $j\in\mathcal{U}$ we have $\phi_{j,t}=\Phi(x_{j},y;x_{j}/\tau,1-r/\tau)$ and we get:
$2B_{k,\ell}\geq d^{2}+\sum_{j\in\mathcal{U}}x_{j}H_{j}\bigl(x_{j}+(a-c_{0}x_{j})(s-x_{j})+2d\Phi(x_{j},y;x_{j}/\tau,1-r/\tau)\bigr)$
and the result follows since $x_{j}\in(0,r]$. ∎
There is one additional complication in the analysis. To take advantage of stochastic processing-time classes, we need a bound for the bonus term accounting for every job $j$ individually, irrespective of how it is clustered with others. We handle this via the following additional approximation:
Proposition 25.
Define parameter $\kappa=0.744$. For a non-leftover cluster $\mathcal{C}^{*}_{k,\ell}$, we have
$B_{k,\ell}\geq c_{1}\sum_{j\in\mathcal{C}^{*}_{k,\ell}}x_{j}(H_{j}-\kappa)\qquad\text{for constant $c_{1}:=0.591909$}.$
Proof.
Let $\lambda,a,r,s,y,d$ be as in Lemma 24, and define $V=\sum_{j\in\mathcal{C}^{*}_{k,\ell}}x_{j}(H_{j}-\kappa)$.
If $\mathcal{C}^{*}_{k,\ell}$ has no truncated job, then $y=d=0$ and $\lambda=r\in[\theta,\tau]$ and $V=s-\kappa r$. So $a=\frac{\mathrm{e}^{1/r}-1}{\mathrm{e}^{1/r}+1}$ and Lemma 24 gives
$\frac{B_{k,\ell}}{V}\geq\inf_{\begin{subarray}{c}r\in[\theta,\tau],x\in(0,r]\\ s\in[r,\pi r]\end{subarray}}\frac{s\bigl(x+(\frac{\mathrm{e}^{1/r}-1}{\mathrm{e}^{1/r}+1}-c_{0}x)(s-x)\bigr)}{2(s-\kappa r)}$
and, as we show in Appendix B, this is at least $c_{1}$.
If $\mathcal{C}^{*}_{k,\ell}$ has a truncated job, then $y\geq\tau-r$, and so $V=s+d-\kappa(r+y)$ and $\lambda=\tau$ and $a=\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}$. Lemma 24 gives:
$\frac{B_{k,\ell}}{V}\geq\inf_{\begin{subarray}{c}r\in[0,\theta],x\in(0,r]\\ y\in[\tau-r,1]\\ s\in[r,\pi r],d\in[y,\pi y]\end{subarray}}\frac{d^{2}+s\bigl(x+(\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}-c_{0}x)(s-x)+2d\Phi(x,y;x/\tau,1-r/\tau)\bigr)}{2(s+d-\kappa(r+y))}$
We show in Appendix B this is at least $0.5921116\geq c_{1}$. ∎
The non-leftover clusters are considered the “baseline” value. We also need to show how the leftover bonus value changes compared to this baseline.
Proposition 26.
For a leftover cluster $\mathcal{C}^{*}_{k,\ell}=\mathcal{C}^{*}_{k,\text{left}}$, define
$R_{k}=\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}x_{j},\qquad S_{k}=\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}x_{j}H_{j},\qquad T_{k}=\frac{S_{k}}{R_{k}}$
Then $R_{k}\in[0,\theta],T_{k}\in[1,\pi]$ and
$B_{k,\ell}\geq R_{k}\cdot\min\{c_{1}(T_{k}-\kappa),c_{2}R_{k}T_{k}^{2}\}\qquad\text{for constant $c_{2}:=0.33965$}$
(If $R_{k}=0$, we may set $T_{k}$ to an arbitrary value in $1,\pi]$ and the result still holds.)
Proof.
The cluster $\mathcal{C}^{*}_{k,\text{left}}$ has no truncated jobs and has $R_{k}\in[0,\theta]$, as otherwise it would be closed after adding its final job. In the language of Lemma 24, we have $r=R_{k},s=S_{k}=R_{k}T_{k}$, and $y=d=0$, and $a\geq\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}\approx 0.67930078\geq 2c_{2}=0.6793$.
Lemma 24 gives $B_{k,\ell}\geq s/2\cdot\inf_{x\in(0,r]}(x+(2c_{2}-c_{0}x)(s-x)\bigr)$. Thus, removing a common factor of $r$ and letting $t=T_{k}$, it suffices to show that
$t/2\cdot(x+(2c_{2}-c_{0}x)(rt-x))\geq\min\{c_{1}(t-\kappa),c_{2}rt^{2}\}.$
This is an algebraic inequality which can be verified to hold for $r\in[0,\theta],t\in[1,\pi],x\in[0,r]$. ∎
We thus can get the following upper bound for the algorithm performance.
Lemma 27.
Define function $f$ and related random variable $D$ by:
$f(r,t)=\max\{0,c_{1}(t-\kappa)-c_{2}rt^{2}\},\qquad\qquad D=\sum_{k}P_{k}^{2}R_{k}f(R_{k},T_{k}),$
We have
${\mathbb{E}}[Z]\leq{\mathbb{E}}[D]+c_{3}Q+L^{2}/2\qquad\text{for constant $c_{3}:=0.8277532$}$
Proof.
By Proposition 26, each leftover cluster $\mathcal{C}^{*}_{k,\text{left}}=\mathcal{C}^{*}_{k,\ell}$ has
$B_{k,\ell}\geq R_{k}\min\{c_{1}(T_{k}-\kappa),c_{2}R_{k}T_{k}^{2}\}=-R_{k}f(R_{k},T_{k})+c_{1}\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}x_{j}(H_{j}-\kappa).$
Putting this together with Proposition 26, we can sum over all clusters in a class $\mathcal{P}^{*}_{k}$ to get:
$\sum_{\ell}B_{k,\ell}$ $\geq\bigl(-R_{k}f(R_{k},T_{k})+c_{1}\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}x_{j}(H_{j}-\kappa)\bigr)+\sum_{\begin{subarray}{c}\text{non-leftover }\mathcal{C}^{*}_{k,\ell}\end{subarray}}c_{1}\sum_{j\in\mathcal{C}^{*}_{k,\ell}}x_{j}(H_{j}-\kappa)$
$=-R_{k}f(R_{k},T_{k})+c_{1}\sum_{j\in\mathcal{P}^{*}_{k}}x_{j}(H_{j}-\kappa)$
Summing over all classes, and noting that $P_{k}=p_{j}/H_{j}$ for $j\in\mathcal{P}^{*}_{k}$, we have
$\sum_{k,\ell}P_{k}^{2}B_{k,\ell}$ $\geq\sum_{k}(-P_{k}^{2}R_{k}f(R_{k},T_{k}))+c_{1}\sum_{k}P_{k}^{2}\sum_{j\in\mathcal{P}^{*}_{k}}x_{j}(H_{j}-\kappa)$
$=-D+c_{1}\sum_{k}\sum_{j\in\mathcal{P}^{*}_{k}}x_{j}(H_{j}-\kappa)p_{j}^{2}/H_{j}^{2}=-D+c_{1}\sum_{j}x_{j}p_{j}^{2}(H_{j}-\kappa)/H_{j}^{2}$
Lemma 23 gives ${\mathbb{E}}[Z\mid P_{\text{offset}}]\leq Q+L^{2}/2-\sum_{k,\ell}P_{k}^{2}B_{k,\ell}$. So we have shown at this point that
${\mathbb{E}}[Z\mid P_{\text{offset}}]$ $\leq Q+L^{2}/2+D+c_{1}\sum_{j}x_{j}p_{j}^{2}(H_{j}-\kappa)/H_{j}^{2}$ (6)
We now integrate over random variable $P_{\text{offset}}$; by Observation 17, for any job $j$ we have
${\mathbb{E}}[(H_{j}-\kappa)/H_{j}^{2}]=\frac{1}{\log\pi}\int_{h=1}^{\pi}\frac{h-\kappa}{h^{3}}\ \mathrm{d}h=\frac{\kappa-2\pi+2\pi^{2}-\kappa\pi^{2}}{2\pi^{2}\log\pi}$
So, in expectation, the term $\sum_{j}x_{j}p_{j}^{2}(H_{j}-\kappa)/H_{j}^{2}$ contributes $\sum_{j}x_{j}p_{j}^{2}\cdot\frac{\kappa-2\pi+2\pi^{2}-\kappa\pi^{2}}{2\pi^{2}\log\pi}$, and thus
${\mathbb{E}}[Z]\leq(1-c_{1}\cdot\frac{\kappa-2\pi+2\pi^{2}-\kappa\pi^{2}}{2\pi^{2}\log\pi})Q+L^{2}/2+D.$
Direct numerical calculation shows that $1-c_{1}\cdot\frac{\kappa-2\pi+2\pi^{2}-\kappa\pi^{2}}{2\pi^{2}\log\pi}\leq c_{3}$. ∎
5.2 The algorithm lower-bound
For this section, we will choose a parameter $\beta:=1.935$, and we define related functions $g,g_{k}$ by
$g(r,t,h)=h\cdot\Bigl(1-\sqrt{1-\frac{\beta(h-\kappa)f(r,t)}{h^{2}(t-\kappa)}}\Bigr),\qquad\qquad g_{k}(h)=g(R_{k},T_{k},h)$
In analyzing functions $f(r,t)$ and $g(r,t,h)$, we implicitly assume throughout that $r\in[0,\theta]$ and $t,h\in[1,\pi]$. It can be checked that in this region $\frac{\beta(h-\kappa)f(r,t)}{h^{2}(t-\kappa)}\leq 1$, so $g$ is well-defined. Likewise, it can be checked that, for any $r,t$, the map $h\mapsto g(r,t,h)$ is non-decreasing and concave-down.
Lemma 28.
We have
$\operatorname{LB}\geq\frac{1}{2}\bigl(Q+\beta D+(L-A)^{2}\bigr)\quad\text{for random variable $A=\sum_{k}P_{k}\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}x_{j}g_{k}(H_{j})$}$
Proof.
Define vector $y\in[0,1]^{\mathcal{J}^{*}}$ as follows: for a job $j\in\mathcal{C}^{*}_{k,\text{left}}$, we set
$y_{j}=\frac{\beta(H_{j}-\kappa)f(R_{k},T_{k})}{H_{j}^{2}(T_{k}-\kappa)}\in[0,1];$
note that $1-\sqrt{1-y_{j}}=g_{k}(H_{j})/H_{j}$. For all other jobs $j$ we set $y_{j}=0$. By Theorem 20, we have:
$2\cdot\operatorname{LB}$ $\geq Q+\sum_{k}\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}\frac{\beta f(R_{k},T_{k})(H_{j}-\kappa)}{H_{j}^{2}(T_{k}-\kappa)}\cdot x_{j}p_{j}^{2}+\Bigl(L-\sum_{k}\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}x_{j}p_{j}\cdot g_{k}(H_{j})/H_{j}\Bigr)^{2}$
$=Q+\beta\sum_{k}P_{k}^{2}\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}x_{j}\frac{f(R_{k},T_{k})(H_{j}-\kappa)}{T_{k}-\kappa}+\Bigl(L-\sum_{k}P_{k}\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}g_{k}(H_{j})x_{j}\Bigr)^{2}$
$=Q+\beta D+(L-A)^{2}.\qed$
We turn to bounding the random variable $A$. This is quite involved and requires a number of intermediate calculations.
Proposition 29.
Define parameter $\delta:=1.3$. We have
$A^{2}\leq\sum_{k}P_{k}^{2}\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}x_{j}g_{k}(H_{j})\Bigl(c_{4}+R_{k}\cdot\frac{T_{k}+\delta}{H_{j}+\delta}\cdot g_{k}(H_{j})\Bigr)$ for constant $c_{4}:=0.035898$.
Proof.
For each class $\mathcal{P}_{k}$, define $A_{k}=\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}x_{j}g_{k}(H_{j})$. We expand the sum $A^{2}$ as:
$A^{2}=\bigl(\sum_{k}P_{k}A_{k}\bigr)^{2}=\sum_{k}P_{k}A_{k}\bigl(P_{k}A_{k}+2\sum_{\ell<k}P_{\ell}A_{\ell}\bigr)=\sum_{k}P_{k}^{2}\bigl(A_{k}^{2}+2A_{k}\sum_{\ell<k}\pi^{\ell-k}A_{\ell}\bigr)$
Since $g_{\ell}(h)$ is a concave-down function, Jensen’s inequality gives
$A_{\ell}\leq\Bigl(\sum_{j\in\mathcal{C}^{*}_{\ell,\text{left}}}x_{j}\Bigr)\cdot g_{\ell}\Bigl(\frac{\sum_{j\in\mathcal{C}^{*}_{\ell,\text{left}}}x_{j}H_{j}}{\sum_{j\in\mathcal{C}^{*}_{\ell,\text{left}}}x_{j}}\Bigr)=R_{\ell}\cdot g(R_{\ell},T_{\ell},T_{\ell})$
Furthermore, since $g(r,t,h)$ is an algebraic function, we can use standard algorithms to calculate
$c_{4}\geq\max_{\begin{subarray}{c}r\in[0,\theta],t\in[1,\pi]\end{subarray}}\frac{2rg(r,t,t)}{\pi-1}$
So $A_{\ell}\leq c_{4}\cdot\frac{\pi-1}{2}$ for each $\ell$; by the geometric series formula, we thus get:
$A^{2}\leq P_{k}^{2}\bigl(A_{k}^{2}+2A_{k}\sum_{\ell<k}\bigl(c_{4}\cdot\tfrac{\pi-1}{2}\bigr)\pi^{\ell-k}\bigr)=\sum_{k}P_{k}^{2}\bigl(A_{k}^{2}+c_{4}A_{k}\bigr)$
To bound the term $A_{k}^{2}$, we use the Cauchy-Schwarz inequality:
$A_{k}^{2}$ $=\Bigl(\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}\sqrt{x_{j}(H_{j}+\delta)}\cdot\frac{g_{k}(H_{j})\sqrt{x_{j}}}{\sqrt{H_{j}+\delta}}\Bigr)^{2}\leq\Bigl(\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}x_{j}(H_{j}+\delta)\Bigr)\cdot\Bigl(\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}\frac{g_{k}(H_{j})^{2}x_{j}}{H_{j}+\delta}\Bigr)$
$=R_{k}(T_{k}+\delta)\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}\frac{x_{j}g_{k}(H_{j})^{2}}{H_{j}+\delta}.\qed$
Lemma 30.
Define parameter $\gamma:=0.00547$. There holds
${\mathbb{E}}[A^{2}]\leq\gamma{\mathbb{E}}[D]+c_{5}Q\qquad\text{for constant $c_{5}:=0.004562$}.$
Proof.
Note that any class $\mathcal{P}^{*}_{k}$ has $\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}x_{j}\cdot\frac{H_{j}+\delta}{T_{k}+\delta}=R_{k}$. Combining the decomposition in Proposition 29 with the formula $D=\sum_{k}P_{k}^{2}R_{k}f(R_{k},T_{k})$, we have:
$A^{2}-\gamma D$ $\leq\sum_{k}P_{k}^{2}\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}x_{j}\Bigl(g_{k}(H_{j})\bigl(c_{4}+\frac{R_{k}(T_{k}+\delta)g_{k}(H_{j})}{H_{j}+\delta}\bigr)-\gamma\frac{H_{j}+\delta}{T_{k}+\delta}f(R_{k},T_{k})\Bigr)$ (7)
Accordingly, let us define function $F:[1,\pi]\rightarrow\mathbb{R}$ by
$F(h)=\max_{\begin{subarray}{c}r\in[0,\theta],t\in[1,\pi]\end{subarray}}\ \ \ g(r,t,h)\bigl(c_{4}+\frac{r(t+\delta)g(r,t,h)}{h+\delta}\bigr)-\frac{\gamma f(r,t)(h+\delta)}{t+\delta}$
Since $R_{k}\in[0,\theta],T_{k}\in[1,\pi]$, Eq. (7) implies that
$A^{2}\leq\gamma D+\sum_{k}P_{k}^{2}\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}x_{j}F(H_{j})=\gamma D+\sum_{k}\sum_{j\in\mathcal{C}^{*}_{k,\text{left}}}x_{j}p_{j}^{2}F(H_{j})/H_{j}^{2}$ (8)
Note that $f(0,0)=g(0,0,h)=0$ for all $h$, and so values $r=t=0$ witness that $F(h)\geq 0$. We can further get an upper bound by summing over all jobs (not just the leftover jobs):
$A^{2}\leq\gamma D+\sum_{j\in\mathcal{J}^{*}}x_{j}p_{j}^{2}F(H_{j})/H_{j}^{2}$
At this point, we take expectations over random variable $P_{\text{offset}}$, getting
${\mathbb{E}}[A^{2}]\leq\gamma{\mathbb{E}}[D]+\sum_{j\in\mathcal{J}^{*}}x_{j}p_{j}^{2}{\mathbb{E}}[F(H_{j})/H_{j}^{2}]$
By Observation 17, each item $j$ has ${\mathbb{E}}[F(H_{j})/H_{j}^{2}]=\frac{1}{\log\pi}\int_{h=1}^{\pi}F(h)/h^{3}\ \mathrm{d}h$, giving us
${\mathbb{E}}[A^{2}]$ $\leq\gamma{\mathbb{E}}[D]+\Bigl(\frac{1}{\log\pi}\int_{h=1}^{\pi}\frac{F(h)}{h^{3}}\ \mathrm{d}h\Bigr)Q$
We show in Appendix B that $c_{5}\geq\frac{1}{\log\pi}\int_{h=1}^{\pi}\frac{F(h)}{h^{3}}\ \mathrm{d}h$, which yields the claimed result. ∎
5.3 Marrying the upper and lower bounds
At this point, we have an upper bound which depends on quantities $Q,D,L$, and we have a lower bound which depends on quantities $Q,D,L,A$. We now can combine these two estimates; critically, we boil down the statistics $Q,D$ into the single parameter:
$q=Q+{\mathbb{E}}[D]/c_{3}$
Observation 31.
There holds ${\mathbb{E}}[Z]\leq c_{3}q+L^{2}/2$
Proof.
Restatement of Lemma 27. ∎
Proposition 32.
There holds ${\mathbb{E}}[(L-A)^{2}]\geq\max\{0,L-c_{6}\sqrt{q}\}^{2}$ for constant $c_{6}:=0.0676$.
Proof.
From Lemma 30, we have ${\mathbb{E}}[A^{2}]\leq\gamma{\mathbb{E}}[D]+c_{5}Q=q\cdot\frac{\gamma{\mathbb{E}}[D]+c_{5}Q}{{\mathbb{E}}[D]/c_{3}+Q}\leq q\cdot\max\{\frac{\gamma}{1/c_{3}},\frac{c_{5}}{1}\}$; it can be checked numerically that this is at most $c_{6}^{2}q$. Now consider random variable $U=A^{2}$. The function $u\mapsto(L-\sqrt{u})^{2}$ is concave-up, so by Jensen’s inequality,
${\mathbb{E}}[(L-A)^{2}]={\mathbb{E}}[(L-\sqrt{U})^{2}]\geq(L-\sqrt{{\mathbb{E}}[U]})^{2}$
The bound ${\mathbb{E}}[A^{2}]={\mathbb{E}}[U]\leq c_{6}^{2}q$ implies that $(L-\sqrt{{\mathbb{E}}[U]})^{2}\geq\max\{0,L-c_{6}\sqrt{q}\}^{2}$. ∎
Proposition 33.
We have
$\operatorname{LB}\geq\frac{\beta c_{3}q+\max\{0,L-c_{6}\sqrt{q}\}^{2}}{\beta c_{3}+1}.$
Proof.
By Corollary 21, we have $\operatorname{LB}\geq Q$. By Lemma 28, we have $\operatorname{LB}\geq\frac{1}{2}\bigl(Q+\beta D+(L-A)^{2}\bigr)$. We can take a convex combination of these two lower bounds to get:
$\operatorname{LB}\geq\alpha Q+(1-\alpha)\cdot\tfrac{1}{2}\bigl(Q+\beta D+(L-A)^{2}\bigr)\qquad\qquad\text{for $\alpha=\frac{\beta c_{3}-1}{\beta c_{3}+1}\approx 0.23$}.$
Taking expectations over $P_{\text{offset}}$ and rearranging, we get:
$\operatorname{LB}$ $\geq Q\cdot\frac{1+\alpha}{2}+\frac{(1-\alpha)\beta}{2}\cdot{\mathbb{E}}[D]+\frac{1-\alpha}{2}\cdot{\mathbb{E}}[(L-A)^{2}]=\frac{\beta c_{3}q+{\mathbb{E}}[(L-A)^{2}]}{\beta c_{3}+1}.$
By Proposition 32, we have ${\mathbb{E}}[(L-A)^{2}]\geq\max\{0,L-c_{6}\sqrt{q}\}^{2}$. ∎
Proposition 34.
The approximation ratio, and the SDP integrality gap, are at most $1.398$.
Proof.
By Corollary 19, we need to bound the ratio ${\mathbb{E}}[Z]/\operatorname{LB}$. Let $v=\sqrt{q}/L$. From Observation 31 and Proposition 33, we have:
$\frac{{\mathbb{E}}[Z]}{\operatorname{LB}}\leq\frac{(\beta c_{3}+1)(c_{3}q+L^{2}/2)}{\beta c_{3}q+\max\{0,L-c_{6}\sqrt{q}\}^{2}}=\frac{(\beta c_{3}+1)(c_{3}v^{2}+1/2)}{\beta c_{3}v^{2}+\max\{0,1-c_{6}v\}^{2}}.$
This is an algebraic function of $v$; its maximum value, over any $v\geq 0$, is at most $1.39798$. ∎
6 Acknowledgments
We thank Aravind Srinivasan for clarifications about the paper [3], Nikhil Bansal for clarifications about the paper [2], and Shi Li for discussions and explanations regarding the time-indexed LP.
Appendix A Proof of Lemma 14
We begin with a few preliminary calculations.
Observation 35.
For $a_{1}\geq a^{\prime}_{1}\geq 1,a_{2}\geq a_{2}^{\prime}\geq 1,b\geq 0$, there holds
$\frac{(a_{1}-1)(a_{2}-1)}{a_{1}a_{2}+b}\geq\frac{(a_{1}^{\prime}-1)(a_{2}^{\prime}-1)}{a^{\prime}_{1}a^{\prime}_{2}+b}.$
Proof.
Can be shown mechanically e.g via decidability of first-order theory of real-closed fields. ∎
Proposition 36.
For $t\geq 4/3$ and $x\in[0,1/t]$, we have $(1-xt)^{1-1/x}\geq\mathrm{e}^{t}(1+x(t^{2}/2-t))$.
Proof.
Consider the function $f(x)=(1-xt)^{1-1/x}$ and let $z=t^{2}/2-t$. Here $e^{t}(1+xz)$ is simply the first-order Taylor expansion of $f$ around $x=0$. Thus, it suffices to show that $f^{\prime\prime}(x)\geq 0$ for $x\in[0,1/t]$. If we denote $y=1-tx\in[0,1]$, we calculate:
$f^{\prime\prime}(x)=\frac{t^{3}y^{-(1-y+t)/(1-y)}}{(1-y)^{4}}\Bigl(t(1-y+\log y)^{2}-(1-y)(1-y^{2}+2y\log y)\Bigr)$
Since $t\geq 4/3$, it thus suffices to show that $4/3\cdot(1-y+\log y)^{2}-(1-y)(1-y^{2}+2y\log y)\geq 0$ holds for $y\in[0,1]$. This is a one-variable inequality which can be shown via straightforward calculus. ∎
We now turn to proving Lemma 14. Rephrasing Theorem 13, we want to show that:
$\frac{((1-x_{1}t)^{1-1/x_{1}}-1)((1-x_{2}t)^{1-1/x_{2}}-1)}{(1-x_{1}t)^{1-1/x_{1}}(1-x_{2}t)^{1-1/x_{2}}+x_{1}t+x_{2}t-1}\geq\frac{\mathrm{e}^{t}-1}{\mathrm{e}^{t}+1}-0.57(x_{1}+x_{2})\max\{0,1/t-0.45\}$
Let $z=t^{2}/2-t$. By combining Observation 35 and Proposition 36, we get
$\frac{((1-x_{1}t)^{1-1/x_{1}}-1)((1-x_{2}t)^{1-1/x_{2}}-1)}{(1-x_{1}t)^{1-1/x_{1}}(1-x_{2}t)^{1-1/x_{2}}+x_{1}t+x_{2}t-1}\geq\frac{(\mathrm{e}^{t}(1+zx_{1})-1)(\mathrm{e}^{t}(1+zx_{2})-1)}{\mathrm{e}^{2t}(1+zx_{1})(1+zx_{2})+x_{1}t+x_{2}t-1}$
Let $d=x_{1}+x_{2},a=\frac{\mathrm{e}^{t}-1}{\mathrm{e}^{t}+1},b=0.57\max\{0,1/t-0.45\}$; to show Lemma 14, it thus suffices to show that
$(\mathrm{e}^{t}(1+zx_{1})-1)(\mathrm{e}^{t}(1+zx_{2})-1)-(a-bd)(\mathrm{e}^{2t}(1+zx_{1})(1+zx_{2})+dt-1)\geq 0$ (9)
We can rearrange the LHS of Eq. (9) as:
$z^{2}x_{1}x_{2}\mathrm{e}^{2t}(1-a+bd)+bd\bigr(\mathrm{e}^{2t}-1+d(t+z\mathrm{e}^{2t})\bigr)+ad(z\mathrm{e}^{t}-t)$
Since $d\in[0,1/t]$ and $a\leq 1$, it suffices to show:
$b\bigr(\mathrm{e}^{2t}-1+\max\{0,1+z\mathrm{e}^{2t}/t\}\bigr)+a\bigl(z\mathrm{e}^{t}-t\bigr)\geq 0$ (10)
where here $a,b,z$ are all functions of the single variable $t$. For $t\geq 2.22$, this is immediate since $z\mathrm{e}^{t}-t\geq 0$. We can divide the remaining search region $[4/3,2.22]$ into strips of width $\varepsilon=1/2000$. Within each strip, we use interval arithmetic in Mathematica to bound the LHS of Eq. (10). Over all regions, it is at least $0.0057>0$.
Appendix B Numerical analysis
Recall that we have the following values for the relevant parameters:
$\pi=3.9$ $c_{0}=0.17556$
$\theta=0.555$ $c_{1}=0.591909$
$\tau=0.604$ $c_{2}=0.33965$
$\beta=1.935$ $c_{3}=0.8277532$
$\delta=1.3$ $c_{4}=0.035898$
$\gamma=0.00547$ $c_{5}=0.004562$
$\kappa=0.744$ $c_{6}=0.0676$
B.1 Calculation of $c_{1}$
Let us first calculate the value
$c_{1}^{\prime}:=\min_{\begin{subarray}{c}r\in[\theta,\tau]\\ s\in[r,\pi r],x\in[0,r]\end{subarray}}\frac{s\cdot\bigl(x+(\frac{\mathrm{e}^{1/r}-1}{\mathrm{e}^{1/r}+1}-c_{0}x)(s-x)\bigr)}{2(s-\kappa r)}$ (11)
Define $t=s/r$. First, if $s\leq 4/3$, then $x+(\frac{\mathrm{e}^{1/r}-1}{\mathrm{e}^{1/r}+1}-c_{0}x)(s-x)$ is an increasing function of $x$. So we can lower-bound it in this range by its value at $x=0$, namely:
$\frac{s^{2}\cdot\frac{\mathrm{e}^{1/r}-1}{\mathrm{e}^{1/r}+1}}{2(s-\kappa r)}=\frac{r(\mathrm{e}^{1/r}-1)}{2\mathrm{e}^{1/r}+2}\cdot\frac{t^{2}}{t-\kappa}\geq\frac{\theta(\mathrm{e}^{1/\theta}-1)}{2\mathrm{e}^{1/\theta}+2}\cdot 4\kappa\approx 0.591909465$...,
where the second inequality comes from straightforward calculus.
Next, we consider the case where $s\geq 4/3$. Since $r\leq\tau$, we can estimate:
$\frac{s\bigl(x+(\frac{\mathrm{e}^{1/r}-1}{\mathrm{e}^{1/r}+1}-c_{0}x)(s-x)\bigr)}{2(s-\kappa)r}\geq\frac{s\bigl(x+(\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}-c_{0}x)(s-x)\bigr)}{2(s-\kappa r)}$ (12)
Since the RHS of Eq. (12) is an algebraic function of $r,s,x$, it can be checked automatically that its mimimum value is $0.6561$... over the domain $r\in[\theta,\tau],s\in[\max\{r,4/3\},\pi r],x\in[0,r]$. Putting the two cases together, we have shown that $c_{1}^{\prime}\geq 0.591909465..$.
We next turn to calculating the value
$c_{1}^{\prime\prime}:=\inf_{\begin{subarray}{c}r\in[0,\theta],x\in(0,r]\\ y\in[\tau-r,1],s\in[r,\pi r]\\ d\in[y,\pi y]\end{subarray}}\frac{d^{2}+s\bigl(x+(\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}-c_{0}x)(s-x)+2d\Phi(x,y;x/\tau,1-r/\tau)}{2(s+d-\kappa(r+y))}$ (13)
where
$\Phi(x,y;x/\tau,1-r/\tau)=\frac{((1-x/\tau)^{1-1/x}-1)((r/\tau)^{1-1/y}-1)}{(1-x/\tau)^{1-1/x}(r/\tau)^{1-1/y}-(r-x)/\tau}.$
From Proposition 36, we note the following bound:
$(1-x/\tau)^{1-1/x}\geq\mathrm{e}^{1/\tau}(1-zx)\qquad\text{for $z=\tau^{-1}-\tau^{-2}/2\approx 0.285$}$
We will divide the search region for $r,y$ into boxes $r_{\min},r_{\max}]\times[y_{\min},y_{\max}]$ of width $\varepsilon=1/1000$. Within a given box (and subject to all other constraints on the variables), we can approximate:
$(1-x/\tau)^{1-1/x}$ $\geq u_{0}(1-zx)$
$(r/\tau)^{1-1/y}$ $\geq(r_{\max}/\tau)^{1-1/y_{\max}}\geq u_{1}$
$(r-x)/\tau$ $\geq(r_{\min}-x)/\tau$
$\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}$ $\geq u_{2}$
for $u_{0},u_{1},u_{2}$ rational numbers within $10^{-7}$ of $\mathrm{e}^{1/\tau},(r_{\max}/\tau)^{1-1/y_{\max}},\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}$ respectively. By Observation 35 and these other approximations, we can lower-bound the expression in Eq. (13) by:
$\frac{d^{2}+s\bigl(x+(u_{2}-c_{0}x)(s-x)+\frac{2d(u_{0}(1-zx)-1)(u_{1}-1)}{u_{0}(1-zx)u_{1}-(r_{\min}-x)/\tau}\bigr)}{2(s+d-\kappa(r_{\min}+y_{\min}))}$ (14)
which is an algebraic function of $x,s,d$ and now no longer depends upon $r$ or $y$. For each such box, we use Mathematica algorithms to minimize over $s,d,x$ subject to constraints $s\in[r_{\min},\pi r_{\max}],d\in[y_{\min},\pi y_{\max}],x\in[0,r_{\max}]$. With $\varepsilon=1/1000$, we calculate a lower bound of $c_{1}^{\prime\prime}\geq 0.5921116$.
B.2 Calculation of $c_{5}$
Recall the functions $f,g,F$ defined as
$f(r,t)=\max\{0,c_{1}(t-\kappa)-c_{2}rt^{2}\},\qquad g(r,t,h)=h\cdot\Bigl(1-\sqrt{1-\frac{\beta(h-\kappa)f(r,t)}{h^{2}(t-\kappa)}}\Bigr)$
$F(h)=\max_{\begin{subarray}{c}r\in[0,\theta]\\ t\in[1,\pi]\end{subarray}}g(r,t,h)(c_{4}+\frac{r(t+\delta)g(r,t,h)}{h+\delta})-\frac{\gamma(h+\delta)f(r,t)}{t+\delta}$
Since $g(r,t,h)$ is a non-decreasing function of $h$, we can get an upper bound within a region $h\in[h_{\min},h_{\max}]$ by:
$F(h)\leq F^{\max}[h_{\min},h_{\max}]:=\max_{\begin{subarray}{c}r\in\mathbb{R},t\in[1,\pi]\end{subarray}}g(r,t,h_{\max})(c_{4}+\frac{r(t+\delta)g(r,t,h_{\max})}{h_{\min}+\delta})-(h_{\min}+\delta)\frac{\gamma f(r,t)}{t+\delta}$
This can be evaluated easily by Mathematica for given $h_{\min},h_{\max}$. Note that the constraint $r\in[0,\theta]$ has been relaxed to $r\in\mathbb{R}$; this is necessary for Mathematica to compute the optimization efficiently. To handle the overall integral $\int_{h=1}^{\pi}F(h)/h^{3}\ \mathrm{d}h$, we divide the integration region $1,\pi]$ into strips of width $\varepsilon=10^{-3}$, and estimate:
$\int_{h=1}^{\pi}\frac{F(h)\ \mathrm{d}h}{h^{3}}\leq\sum_{i=0}^{\lceil(\pi-1)/\varepsilon\rceil-1}\int_{h=1+i\varepsilon}^{1+(i+1)\varepsilon}\frac{F^{\max}[1+i\varepsilon,1+(i+1)\varepsilon]\ \mathrm{d}h}{h^{3}}$
$\qquad=\sum_{i=0}^{\lceil(\pi-1)/\varepsilon\rceil-1}F^{\max}[1+i\varepsilon,1+(i+1)\varepsilon]\cdot\Bigl(\frac{1}{2(i\varepsilon+1)^{2}}-\frac{1}{2((i+1)\varepsilon+1)^{2}}\Bigr)$
which is calculated to be at most $c_{5}$.